### v1

# DBellQuant: Breaking the Bell with Double-Bell Transformation for LLMs Post Training Binarization

Abstract
Large language models (LLMs) demonstrate remarkable performance but face substantial computational and memory challenges that limit their practical deployment. Quantization has emerged as a promising solution; however, its effectiveness is often limited by quantization errors arising from weight distributions that are not quantization-friendly and the presence of activation outliers. To address these challenges, we introduce DBellQuant, an innovative post-training quantization (PTQ) framework that achieves nearly 1-bit weight compression and 6-bit activation quantization with minimal performance degradation. DBellQuant uses Learnable Transformation for Dual-Bell (LTDB) algorithm, which transforms single-bell weight distributions into dual-bell forms to reduce binarization errors and applies inverse transformations to smooth activations. DBellQuant sets a new state-of-the-art by preserving superior model performance under aggressive weight and activation quantization. For example, on the Wikitext2 dataset, DBellQuant achieves a perplexity of 14.39 on LLaMA2-13B with 6-bit activation quantization, significantly outperforming BiLLM’s 21.35 without activation quantization, underscoring its potential in compressing LLMs for real-world applications.
1 Introduction
In recent years, the rapid advancement of large language models (LLMs) demonstrate exceptional performance in a variety of complex tasks that involve natural language understanding and generation achiam2023gpt ; dubey2024llama . However, these models often comprise hundreds of billions of parameters, posing significant challenges for their deployment in real-world applications because of the substantial computational and memory requirements (e.g. a 70B model requires around 150GB GPU memory), resulting in huge operational costs and unacceptable inference latency.
In this context, quantization, as an effective model compression technique, has garnered significant attention. Due to the sparsity of information density in the weights of LLMs wang2024bitnet ; yu2024super , weight binarization has emerged as a promising quantization scheme. Methods such as PB-LLM shang2023pb and BiLLM huang2024billm have significantly reduced the errors introduced by binarization by applying finer-grained processing to critical weights. Despite its theoretical advantages, the application of quantization in compressing LLMs continues to face numerous challenges egashira2024exploiting . For instance, systematic outliers in activation values can lead to substantial quantization errors, negatively impacting model accuracy sun2024massive . Although recent studies xiao2023smoothquant ; shao2023omniquant have proposed mitigating such errors by redistributing the scaling factors between weights and activation values. Other research directions ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm focus on leveraging Hadamard transformations to effectively eliminate outlier features. However, no binarization methods have taken care of simultaneous smoothing activation outliers. Consequently, the demand of higher bit-width activations during computation leads to significant computational overhead. BitNet a4.8 wang2024bitnet addresses this issue by employing quantization-aware training (QAT) to achieve a 1-bit LLM with 4-bit activations. Nevertheless, the QAT approach requires extensive computational resources and prolonged training time. Therefore, developing post-training quantization (PTQ) methods for binary compression that minimize performance loss while simultaneously smoothing activation and thus reducing activation bit-width remains a critical and challenging task.
Figure 1: Performance on Wikitext2 dataset. DBellQunat outperforms weight-only quantization method under 8-bit activation setting.
To this end, we begin by revisiting the distribution characteristics of activations and weights in LLMs (Fig. 2(a)). We observe that the unimodal nature of weight distributions leads to substantial quantization errors, particularly in the case of low 1-bit quantization. Ideally, a dual-bell-shaped weight distribution (Fig. 2(b)) can effectively reduce binarization errors. This motivates a key question: Is it possible to transform weights into a dual-bell shape distribution while simultaneously addressing activation outliers to facilitate both activation quantization and weight binarization?
Building on these observations, we propose DBellQuant, a novel weight-activation quantization framework for efficient post-training quantization (PTQ). DBellQuant enables activation quantization while achieving near 1-bit weight compression with minimal accuracy loss. At the core of DBellQuant is a learnable transformation matrix that maps the weight distribution into a dual-bell form, while its inverse is applied to the input activations to smooth them. We begin by analyzing the characteristics of weight distributions suitable to binarization and theoretically derive the feasibility of applying a dual-bell transformation. Based on this, we then develop an efficient and lightweight algorithm, Learnable Transformation for Dual-Bell (LTDB). LTDB initializes the transformation matrix using an activation-aware strategy and optimizes it via a custom objective function that encourages the weights to cluster around two centers. This drives the formation of a symmetric dual-peak distribution, and an early stopping strategy is employed to ensure stable and efficient optimization. Through this equivalence-preserving transformation, weights originally exhibiting unimodal distributions– challenging for binarization– are mapped into near-symmetric dual-bell distributions (Fig. 2(b)), significantly reducing binarization error. Simultaneously, the inverse of the learned transformation is applied to the input activations, effectively scaling down outliers and smoothing the distribution. This makes the activations more amenable to quantization without altering the model’s output (Fig. 2(b)).
Experimental results demonstrate that the equivalent transformation strategy of DBellQuant significantly reduces the loss caused by weight binarization. For the first time under PTQ conditions, it achieves near 1-bit weight compression while simultaneously compressing activations to 6 bits.Across various LLMs and evaluation metrics, DBellQuant consistently achieves state-of-the-art results. For instance, on the Wikitext2 benchmark, we achieves a perplexity of 14.39 on LLaMA2-13B using only 6-bit activations (Fig. 1), significantly surpassing the performance of BiLLM, a method that only quantizes weights, which achieves a perplexity of 21.35.
Figure 2: (a) Before applying DBellQuant, activations exhibit significant outliers, making quantization challenging, while the single-bell-shaped weight distribution hinders binarization. (b) After applying DBellQuant, activations are smoothed with substantially fewer outliers, facilitating easier quantization. Weight distribution is transformed to dual-bell form, which is more conducive to binarization.
2 Related Work
Quantization for Large Language Models
The massive parameter size of LLMs poses significant challenges in terms of memory consumption and computational efficiency. Therefore, quantization is crucial to compress these models, reducing resource requirements while preserving performance for practical deployment. LLMs quantization have introduced a variety of innovative techniques to enhance efficiency while maintaining accuracy. Works like GPTQ frantar2022gptq and OBQ frantar2023optimalbraincompressionframework minimizes reconstruction error by adjusting the remaining unquantized parameters in the block to compensate for the accuracy loss caused by quantization. LLM.int8()dettmers2022gpt3 and ZeroQuantyao2022zeroquant improve quantization accuracy by introducing additional grouping labels for customized quantization blocks. Other works like SmoothQuantxiao2023smoothquant and OmniQuant shao2023omniquant addresses activation outliers by redistributing scaling factors between weights and activations, migrating the quantization difficulty from activation to weights. Additionally, recent approaches leverage Hadamard transformations to suppress activation outliers ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm , while incoherence processing has been proposed for effective low-bit quantization chee2024quip2bitquantizationlarge ; tseng2024quipbetterllmquantization . Collectively, these advancements demonstrate that quantization techniques can be successfully scaled to multi-billion-parameter models, achieving substantial reductions in memory consumption and inference latency without compromising model performance.
Binary Quantization
Binary quantization, an extreme low-bit quantization technique that reduces model weights and activations to binary values (e.g., -1 and +1 or 0 and 1), has gained significant attention for its ability to drastically cut memory usage and computational complexity, making it ideal for resource-constrained devices and efficient deployment of large-scale models. However, applying binary quantization to LLMs presents substantial challenges due to their sensitivity to precision loss, particularly in attention mechanisms and large embedding layers. BinaryBERT bai2020binarybert explored binary quantization for BERT, proposing selective preservation of critical weights in higher precision to mitigate performance degradation. In another direction, PB-LLM shang2023pb introduced a partially-binarized approach for LLMs, retaining a small fraction of salient weights in higher precision while binarizing the rest, enabling extreme low-bit quantization without sacrificing linguistic reasoning capabilities. Recent advancements include structural binarization techniques that leverage novel sparsity forms and standardized importance metrics to selectively binarize and sparsify LLM weights dong2024stbllmbreaking1bitbarrier , as well as strategies like alternating refined binarization and column-group bitmap methods to effectively reduce quantization error and address column deviations li2024arbllmalternatingrefinedbinarizations . These innovations collectively advance the feasibility of binary quantization for LLMs, pushing the boundaries of efficiency without compromising performance.
Figure 3: DBellQuant Framework Overview: (a)First, we can see that the origin weight distribution is single-bell. (b)We utilize Activation-aware initialization to generate origin transformation matrix. (c)We employ the LTDB algorithm for iterative training of the transformation matrix, applying the proposed Dual-Transformation Loss in two ways: for training and as the termination criterion for the training process. (d)The weight distribution after transformation will be double-bell.
3 Method
We begin by exploring the process of binarization, analyzing and theoretically proving the weight distributions suitable for binarization in Sec. 3.1. Based on our analysis, we propose the Learnable Transformation for Dual-Bell (LTDB) algorithm in Sec. 3.2. After investigating the potential of utilizing a learnable transformation matrix to achieve the objective, we redesigned an efficient learnable transformation along with a reasonable activation-aware initialization method, taking into account training difficulty and task complexity. An overview of the algorithm is included in Fig. 3.
3.1 Binarization-Friendly Weight Redistribution
By utilizing the sign function, binarization can convert weights in LLMs into binary values. The per-channel binarization and de-binarization process is as follows:
$\small\beta=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{W_{i,j}},\quad\boldsymbol{\widetilde{W}}=\text{Sign}(\boldsymbol{W}-\beta),\quad\alpha=\frac{1}{n}\sum_{i=1}^{n}|\boldsymbol{W_{i,j}}-\beta_{j}|$ (1)
$\text{Sign}(\boldsymbol{W_{i,j}})=\begin{cases}+1,&\text{if }\boldsymbol{W_{i,j}}>0,\\ -1,&\text{if }\boldsymbol{W_{i,j}}\leq 0,\end{cases},\quad\boldsymbol{W_{deq}}=\boldsymbol{\widetilde{W}}\cdot\alpha+\beta$ (2)
where $\beta$ is the shifting factor and $\alpha$ is the scaling factor for binarization. Previous studies huang2024slim have shown that neural network weights exhibit structured distributions along the channel dimension, with certain channels being more salient. The overall weight distribution typically follows a quasi-Gaussian pattern dettmers2022llmint88bitmatrixmultiplication , as does the channel-wise distribution (Fig. 6). Binarizing such weight matrices introduces significant quantization errors, which can severely degrade model performance.
In LLM binarization, a dual-bell distribution is theoretically more advantageous than a single-bell distribution due to its natural separation into two distinct clusters, which aligns well with binary quantization levels (e.g., -1 and 1), thereby minimizing quantization error. In contrast, single-bell distributions, concentrated around a single peak, often cause significant overlap when mapped to binary values, reducing representation accuracy (see Appendix A.4 for detailed analysis). However, LLM weight distributions typically exhibit single-bell characteristics, and conventional PTQ methods fail to effectively transform them for binarization. While QAT can reshape weight distributions into a dual-bell form through its learning objectives wang2023bitnetscaling1bittransformers , it requires substantial computational resources and prolonged training. To address this, we propose a more efficient PTQ method that rapidly converts single-bell distributions into dual-bell ones, optimizing binary quantization without the need for resource-intensive retraining.
3.2 Learnable Transformation for Dual-Bell Quantization
Learnable Transformation with Auxilary Matrix
As mentioned before, double-bell distributions are advantageous for binarization. However, the key question lies in how to transform a weight matrix that originally follows a single-bell distribution into a double-bell one and ensures the computational results remain unchanged. In this section, we first explore the feasibility of achieving such a transformation through the application of an auxiliary matrix:
Theorem 1.
Let $\boldsymbol{W}\in\mathbb{R}^{n\times m}$ be a weight matrix where each channel $\boldsymbol{w}_{i}$ (for $i\in\{1,2,\dots,n\}$) is sampled from a single-bell Gaussian distribution $\boldsymbol{w}_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. There exists a learnable matrix $T\in\mathbb{R}^{m\times m}$, such that the channels of the transformed matrix $\boldsymbol{W}^{\prime}=\boldsymbol{W}T$ follow a double-bell distribution, specifically a mixture of two Gaussians:
$\boldsymbol{w}_{i}^{\prime}\sim\pi\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(\mu_{2},\sigma_{2}^{2}),$
where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are parameters of the doubel-bell distribution. More detailed proof is shown in Appendix. A.3.
Theorem 1 demonstrates that transforming weight distributions from single-bell to double-bell can be achieved by introducing an auxiliary matrix $T$. However, this approach presents several significant challenges. First, in LLMs, weight matrices typically have extremely high dimensionalities, such as $(4096,4096)$, meaning that the auxiliary matrix $T$ would also be of similarly large dimensions, making it computationally expensive and difficult to learn. Second, to maintain computational consistency, it is necessary to simultaneously apply $T^{-1}$ to the activations, which raises a critical issue regarding the invertibility of $T$. Ensuring strict invertibility introduces additional constraints and complex design steps, further complicating the process. Third, even if $T$ is strictly invertible, it remains uncertain whether this design effectively facilitates activation quantization, as there are no explicit mechanisms in the current approach to optimize activation quantization. These limitations highlight the need for a more efficient and robust design to address the computational and practical challenges associated with auxiliary matrix-based transformations.
Learnable Equivalent Transformation
To address these challenges, we propose a simpler and more efficient method for achieving the transformation. In this approach, the matrix $T\in\mathbb{R}^{1\times C_{\text{in}}}$ to be learned is reduced to a $1\times C_{\text{in}}$ matrix. Compared to the matrix introduced above, this significantly reduces the dimensionality of $T$ to compared to its original size, making it substantially easier to learn.
Furthermore, this approach allows for straightforward transformations to ensure computational consistency without introducing additional complexity as follows:
$\small\textbf{Y}=\textbf{X}*\boldsymbol{W}=\textbf{X}*(T^{-1}*T)*\boldsymbol{W}=\left(\textbf{X}\odot T^{-1}\right)*\left(T\odot\boldsymbol{W}\right)$ (3)
where $X\in\mathbb{R}^{N\times C_{\text{in}}}$ is the input matrix, $N$ is the token length and $C_{\text{in}}$ is the input channel size. $w\in\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$ is the weight matrix, where $C_{\text{out}}$ is the output channel size. $\odot$ denotes elementwise multiplication.Moreover, this equivalent transformation matrix $T$ will be directly fused into the LayerNorm weights and the corresponding linear weights, without introducing any additional parameters. However, directly solving this matrix is highly challenging. To address this, we propose a learnable approach to train and derive the matrix effectively.
Here we introduce the way to initialize the learnable transformation matrix $T$ using the following equation:
$\small T_{j}=\frac{\max(|\mathbf{X}_{j}|)^{\epsilon}}{\max(|\boldsymbol{W}_{j}|)^{1-\epsilon}}$ (4)
where $\epsilon$ is a hyperparameter. This initialization strategy provides significant advantages for both weight binarization and activation smoothing. For weight quantization, specifically, when $\max(|\boldsymbol{W}_{j}|)$ is particularly small, it indicates that the absolute value of weights are relatively small, resulting in a large value of $\frac{1}{\max(|\boldsymbol{W}_{j}|)}$ which corresponds to scaling up these smaller weights. Conversely, when $\max(|\boldsymbol{W}_{j}|)$ is particularly large, it reflects larger absolute value weights, which lead to a smaller value of $\frac{1}{\max(|\boldsymbol{W}_{j}|)}$, which will scale down the larger weights. All values can be shifted closer to two central points through these two processes, and it will reduce the quantization error shown in Appendix. A.4. Regarding activation quantization, the initialization explicitly accounts for outliers in the activation matrix, making it inherently activation-aware. This ensures that even without further optimization of activation quantization during subsequent training, it supports near 1-bit weight quantization while effectively reducing activations to a low bit-width.
3.3 Dual-Transformation Optimizing Objectives
Dual-Target Minimum Deviation Loss
Our learning objective is to encourage all weight values to move closer to the two mean centers calculated by Eq. 2, ultimately forming a doubel-bell distribution with these two values as its peaks. During the binarization process, we denote these two points as $m_{1}$ and $m_{2}$ respectively. The simple way to set loss function is as follows:
$\small\mathcal{L}_{\text{DTMD}}=\frac{\lambda_{\text{DTMD}}}{n}\sum_{i=1}^{n}\min\left(\left|\boldsymbol{W}*{T_{i}}-m_{1,i}\right|,\left|\boldsymbol{W}*{T_{i}}-m_{2,i}\right|\right)$ (5)
where $\lambda_{\text{DTMD}}$ represents the coefficient of $\mathcal{L}_{\text{DTMD}}$. However, employing this type of loss function to train the transformation matrix introduces an issue. Specifically, we observed that the transformation matrix tends to shrink progressively during training, contrary to the intended effect of scaling up the originally smaller absolute values. This unintended behavior results in a significant problem: it effectively shifts the quantization challenge from the weights to the activations.
Dual-Target Normalized Proportional Loss
As discussed before, DTMD alone is not enough to train a better weights distribution for binarization. So we introduce a new loss function as follows:
$\small\mathcal{L}_{\text{DTNP}}=\frac{\lambda_{DTNP}}{n}\sum_{i=1}^{n}\begin{cases}\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|}{\left|\text{m}_{1,i}\right|},&\text{if }\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|<\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|\\[10.0pt] \frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|}{\left|\text{m}_{2,i}\right|},&\text{otherwise.}\end{cases}$ (6)
Figure 4: Dual-Target Minimum Deviation Loss value over iterations across different layers.
where $\lambda_{DTNP}$ represents the coefficient of $\text{Loss}_{\text{rel}}$. By leveraging the dual-target normalized proportional objective, the transformation matrix can be effectively trained to meet the desired behavior, scaling down larger absolute values and scaling up smaller absolute values to approach a double-bell-shaped distribution. Since our target values have been transformed into $\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{i}\right|}{\left|\text{m}_{i}\right|}$ , the final convergence values might not align with our desired results. Furthermore, we propose an early stopping mechanism to prevent the function from converging to an undesired solution that deviates from our intended objective. We observe that by using this loss to train, DTMD drops quickly first and then slowly grow as shown in Fig. 4. Therefore, we introduce an early stop mechanism and DTMD is utilized as a condition for stopping the training.
3.4 Impact of the Inverse of Learnable Transformation Matrix on Activation Smoothing
Figure 5: Visualization of distribution of values in $T$ and $T^{-1}$ of Llama2-7B.
Through the use of DTNP for training, we understand that the transformation matrix $T$ drives all values in the weights closer to the two mean centers calculated by Eq. 2. Prior research has shown that the distribution of weights tends to approximate a quasi-Gaussian distribution, with the majority of values being extremely small and close to zero, while only a very small fraction exhibit relatively large absolute values dettmers2022llmint88bitmatrixmultiplication . Theoretically, this implies that $T$ will contain many values greater than 1 to amplify the numerous near-zero absolute values, bringing them closer to the mean centers. At the same time, very few values of $T$ will be less than 1 to reduce the relatively rare large absolute values, aligning them similarly with the mean centers. In fact, when we visualize $T$ after training as shown in Fig. 5, we observe that less than 5% of its values are below 1. Consequently, the corresponding $T^{-1}$, which is multiplied with activations, has over 95% of its values below 1. This significantly reduces the quantization range of activations and suppresses the magnitude of outlier values within the activations. Visualization results can be seen in Appendix. A.6. As a result, this approach is particularly effective in further facilitating activation quantization.
3.5 Algorithm
The process of Learnable Transformation for Dual-Bell Transformation Algorithm is shown in Algorithm. 1, which adjusts the full-precision weight matrix $W$ using an activation-aware initialization transformation matrix $T$ over $N$ epochs. The algorithm iteratively minimizes dual-target loss functions to guide $W$ toward a dual-bell distribution while employing an early stopping mechanism based on the DTMD.
Algorithm 1 Learnable Transformation for Dual-Bell Transformation (LTDB)
1:function LTDB($\boldsymbol{W},\mathbf{T},N$)
2:     Input: $\boldsymbol{W}\in\mathbb{R}^{n\times m}$ a full-precision weight matrix.
3:     Input: $\mathbf{T}\in\mathbb{R}^{1\times m}$ an activation-aware initialization transformation matrix.
4:     Input: $N$ the total number of epochs.
5:     Output: $\widetilde{\boldsymbol{W}}\in\mathbb{R}^{n\times m}$ the transformed weight matrix.
6:     for $\text{iter}=1,2,\dots,N$ do
7:         $\widetilde{\boldsymbol{W}}\leftarrow\mathbf{T}\odot\boldsymbol{W}$triangleright Perform element-wise multiplication
8:         $\mathcal{L}_{\text{DTMD}},\mathcal{L}_{\text{DTNP}}\leftarrow\text{LossFunc}(\widetilde{\boldsymbol{W}})$triangleright Compute the two dual-target losses for $\widetilde{\boldsymbol{W}}$
9:         $\mathcal{L}_{\text{DTNP}}.backward()$triangleright Use Dual-Target Normalized Proportional Loss for training
10:         if $\mathcal{L}_{\text{DTMD}}>\mathcal{L}_{\text{DTMD-Minimum}}$ then
11:              break $\triangleright$ Stop training based on Dual-Target Minimum Deviation Loss
12:         end if
13:         $\mathcal{L}_{\text{DTMD-Minimum}}\leftarrow\mathcal{L}_{\text{DTMD}}$
14:     end for
15:     return $\widetilde{\boldsymbol{W}}$
16:end function
4 Experiments
4.1 Settings
All experimental procedures were executed utilizing the PyTorch paszke2019huggingface framework in conjunction with the Huggingface library paszke2019huggingface . Models with parameters smaller than 8B are running on a single NVIDIA A30 GPU equipped with 24GB of memory, others are running on a single NVIDIA A100 GPU equipped with 80GB of memory. Consistent with methodologies outlined by Frantar et al. frantar2023gptq and Huang et al. huang2024billm , a calibration dataset comprising 128 samples sourced from the C4 collection raffel2020exploring was employed.
Models and Datasets
Comprehensive evaluations were carried out across several large language model families, including LLaMA, LLaMA-2, and LLaMA-3 touvron2023llama and the OPT series zhang2022opt . The efficacy of the developed DBellQuant was assessed by calculating the perplexity of the models’ generated text on standard benchmarks: WikiText2 merity2017pointer , and a subset of the C4 data raffel2020exploring . Furthermore, the models’ performance was evaluated based on accuracy across seven zero-shot question-answering tasks: ARC-c clark2018think , ARC-e clark2018think , Hellaswag zellers2019hellaswag , PIQA bisk2020piqa , and Winogrande sakaguchi2020winogrande .
Comparison Methods
The primary benchmark for comparison for our DBellQuant approach is BiLLM huang2024billm , which represents the current baseline PTQ technique for binary large language models. Additionally, we include other contemporary PTQ algorithms in our comparison, namely Round-to-Nearest (RTN), GPTQ frantar2023gptq , PB-LLM shang2023pb and the current state-of-the-art PTQ technique ARB-LLM li2024arbllmalternatingrefinedbinarizations .
[…]
Method Activation Bits OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-1-7B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
Full Precision 16 14.62 12.47 10.86 5.68 05.47 04.88 3.32
RTN 16 17165.72 36516.69 11550.91 168388.00 157058.34 47902.32 160389.91
GPTQ 16 14844.73 14114.58 10622.81 267001.72 115905.67 9387.80 14219.35
PB-LLM 16 265.52 124.35 105.16 102.36 69.20 151.09 28.37
BiLLM 16 69.97 49.55 35.36 35.04 32.48 21.35 13.32
[…]
BiLLM 8 88.95 68.60 166.46 40.13 33.23 22.55 14.72
[…]
BiLLM 6 9537 18405.85 28123.58 71.65 42.41 30.20 18.65
[…]
4.2 Main Results
[…]
Table 2: Accuracy of PIQA, ARC-e, ARC-c, HellaSwag, Winogrande and average accuracy of all datasets with BiLLM and our methods on OPT and LLaMA family.
Model Method Activation Bits PIQA ARC-e Arc-c HellaSwag Winogrande Avg.
- 16 76.33 65.61 30.55 50.51 65.35 57.67
BiLLM 16 59.63 36.83 17.06 30.14 51.30 38.99
DBellQuant 16 70.29 55.76 24.72 37.81 58.71 49.46
[…]
DBellQuant 8 69.10 54.63 24.91 38.19 57.38 48.84
[…]
DBellQuant 6 68.12 52.44 23.38 37.04 57.30 47.65
- 16 78.40 67.34 38.14 56.45 67.01 61.46
BiLLM 16 61.92 38.93 21.58 32.78 53.67 41.77
DBellQuant 16 67.74 49.37 24.23 39.55 58.80 47.94
[…]
DBellQuant 8 67.41 47.31 26.19 38.89 58.80 47.72
[…]
DBellQuant 6 65.29 46.46 25.77 37.28 54.85 45.94
- 16 78.40 69.28 40.02 56.69 67.25 62.32
BiLLM 16 60.34 36.87 21.59 30.24 51.62 40.13
DBellQuant 16 63.98 42.85 23.89 34.82 56.27 44.36
[…]
DBellQuant 8 62.56 42.42 23.03 34.44 54.38 43.37
[…]
DBellQuant 6 61.10 37.5 22.18 31.94 53.85 41.32
- 16 79.65 80.09 50.51 60.18 72.77 68.64
BiLLM 16 57.51 33.75 18.52 31.63 53.12 38.90
DBellQuant 16 62.35 44.11 19.97 33.19 55.96 43.12
[…]
DBellQuant 8 61.70 40.95 18.40 32.95 55.09 41.82
[…]
DBellQuant 6 58.00 37.63 18.77 31.83 51.92 39.64
4.3 Ablation Experiments
Optimization Objective
In the definition of DTNP, we employed the L1 loss as described in Eq. 6. To evaluate the effectiveness of the L1 loss, we tested by replacing it with L2 loss and do the same experiments. We observed that the overall performance was inferior to that achieved by L1 loss. Therefore, we selected L1 loss as the loss function for DTNP. Additionally, during the activation-aware initialization process, the parameter $\epsilon$ is introduced as a hyperparameter. We conducted experiments with different values of $\epsilon$, and the results demonstrated that all tested values of $\epsilon$ consistently outperformed prior algorithms. For LLaMA-2-7B, setting $\epsilon$ to 0.85 yielded the best performance. Results are shown in Tab. 4.
Impact of Block Size
We investigated the impact of block size on the quantization performance of DBellQuant by experimenting with block sizes ranging from 64 to 256 columns. The results are shown in Tab. 3. Consistent with other PTQ methods, smaller block sizes yield lower perplexity, but it will increase the diversity of quantization results also increase the weighting overhead. In all experiments shown before we set the block size to 128 because we think it can better balance the bit-width and quantization effect. Furthermore, our method consistently outperforms the current baseline approaches across various block sizes, demonstrating its robustness and generalizability under different configurations.
Figure 6: Top: Visualization of single-bell weights distribution from different blocks of different layers before applying DBellQuant. Bottom: Visualization of dual-bell weights distribution from different blocks of different layers after applying DBellQuant.
Activation Bit-width Comparisons
We compare the results across different models with different activation bit-widths shown in Fig. 7. Specifically, we observe that when activations are quantized to 8 bits, the model’s perplexity remains nearly unchanged and, in some cases, even decreases. This demonstrates that our approach alleviates the challenges associated with activation quantization. Furthermore, we found that even under lower-bit quantization, such as 6-bit, our model is able to largely maintain its original performance, with only a minimal increase in perplexity. It is noteworthy that for large-scale models such as LLaMA-2 13B and 70B, the perplexity degradation is almost negligible, highlighting the effectiveness and efficiency of our approach in models with substantial parameters.
Method Avtivation Bits Block Size WikiText2 $\downarrow$ C4 $\downarrow$
BiLLM 16 64 20.12 24.46
DBellQuant 8 64 13.67 16.99
DBellQuant 6 64 15.65 19.71
[…]
DBellQuant 8 128 18.65 23.02
DBellQuant 6 128 21.69 25.91
[…]
DBellQuant 8 256 22.34 23.37
DBellQuant 6 256 24.68 28.52
Table 3: Performance of different block size.
Figure 7: Performance of different activation bit-widths.
Method Loss Function Avtivation Bits WikiText2 $\downarrow$ C4 $\downarrow$
DBellQuant L2 8 18.90 24.11
DBellQuant L1 8 18.65 23.02
[…]
DBellQuant L1 6 21.69 25.91
(a) Performance of different loss function.
Method $\epsilon$ WikiText2 $\downarrow$ C4 $\downarrow$
DBellQuant 0.75 19.57 22.92
DBellQuant 0.8 18.66 22.81
DBellQuant 0.85 17.91 21.83
DBellQuant 0.9 18.29 22.52
(b) Performance of different $\epsilon$
Table 4: Ablation study on LLaMA-2-7B,results are measured by perplexity.
5 Conclusion
In this work, we propose DBellQuant, an efficient PTQ method that enables simultaneous weight binarization and activation quantization for LLMs. By analyzing weight distributions conducive to binarization, we design a novel and efficient Learnable Transformation for Dual-Bell algorithm. Specifically, we introduce two customized loss functions and an early stopping mechanism to achieve the desired dual-bell transformation. Furthermore, we analyze and demonstrate how this transformation improves activation quantization, allowing us to achieve near 1-bit weight compression and 6-bit activation quantization with minimal performance degradation—marking the first time such results have been achieved under a PTQ framework. Experiments on open-source LLM families show that DBellQuant significantly advances the performance frontier of SOTA binary PTQ methods.
Appendix A Appendix
A.1 Limitations
Currently, our work only supports quantizing activations to 6 bits. When attempting to quantize activations to lower bit-widths, the model collapses, resulting in a significant drop in performance. However, some recent studies have demonstrated the ability to quantize activations to 4 bits while maintaining competitive model performance. Inspired by these advancements, we aim to adopt similar approaches to further optimize our models. Specifically, we seek to not only binarize weights but also quantize activations to even lower bit-widths, enabling easier deployment and faster computation.
A.2 Broader Impacts
Our work demonstrates the feasibility of simultaneously binarizing weights and quantizing activations to 6 bits in LLMs while maintaining competitive performance. This approach significantly reduces the computational and memory overhead associated with LLM deployment, making them more accessible for resource-constrained environments. By enabling efficient inference, our method contributes to the democratization of advanced AI technologies, reducing the environmental impact of large-scale model deployment. Furthermore, it opens new avenues for research in ultra-low-bit quantization, fostering innovation in model efficiency and scalability.
A.3 Proof for Theorem. 1
Proof.
Problem Setup: By assumption, the rows of the original weight matrix $\mathbf{W}$ are sampled independently from Gaussian distributions:
$\mathbf{w}_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2}),\quad\text{where }\mu_{i}\in\mathbb{R}\text{ and }\sigma_{i}>0\text{ for all }i.$
We aim to learn a transformation matrix $\mathbf{T}\in\mathbb{R}^{m\times m}$ such that the rows of the resulting matrix $\mathbf{W}^{\prime}=\mathbf{W}\mathbf{T}$ follow a bimodal distribution.
Learnable Transformation Definition: The transformed matrix is defined as:
$\mathbf{W}^{\prime}=\mathbf{W}\mathbf{T},$
where $\mathbf{T}$ is a learnable matrix that modulates the distribution of each row $\mathbf{w}_{i}^{\prime}$ of $\mathbf{W}^{\prime}$. Since the rows of $\mathbf{W}$ are Gaussian-distributed, the linear transformation by $\mathbf{T}$ initially results in a new Gaussian distribution for each row:
$\mathbf{w}_{i}^{\prime}\sim\mathcal{N}(\mu_{i}^{\prime},\sigma_{i}^{\prime 2}),$
where $\mu_{i}^{\prime}=\mu_{i}\mathbf{T}$ and $\sigma_{i}^{\prime 2}=\mathbf{T}^{\top}\Sigma_{i}\mathbf{T}$, with $\Sigma_{i}=\mathrm{diag}(\sigma_{i}^{2})$ being the covariance of $\mathbf{w}_{i}$.
Inducing a Bimodal Distribution: To map the Gaussian-distributed rows $\mathbf{w}_{i}^{\prime}$ into a bimodal distribution, we note that a bimodal distribution can be expressed as a Gaussian mixture model:
$g(x)=\pi\mathcal{N}(x;\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(x;\mu_{2},\sigma_{2}^{2}),$
where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are the parameters of the mixture components. To achieve this, $\mathbf{T}$ is learned to ensure that the linear transformation $\mathbf{W}\mathbf{T}$ reshapes the original Gaussian distribution into a mixture of two Gaussians.
Parameter Optimization: The learnable matrix $\mathbf{T}$ is optimized using a loss function $L$ that minimizes the Kullback-Leibler (KL) divergence between the empirical distribution of the rows of $\mathbf{W}^{\prime}$ and the target bimodal distribution:
$L=\mathrm{KL}\left(p(\mathbf{w}_{i}^{\prime})\,\|\,\pi\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(\mu_{2},\sigma_{2}^{2})\right).$
The optimization process adjusts the entries of $\mathbf{T}$ to align the transformed rows $\mathbf{w}_{i}^{\prime}$ with the desired bimodal distribution.
Conclusion: The existence of such a learnable matrix $\mathbf{T}$ ensures that the rows of the transformed matrix $\mathbf{W}^{\prime}=\mathbf{W}\mathbf{T}$ can follow a bimodal distribution. This completes the proof. ∎
A.4 Analysis of the reasons doubel-bell distribution more suitable for binarization compared to a single-bell distribution.
Directly proving that a dual-bell distribution is more suitable for binarization compared to a single-bell distribution can be challenging, as it requires setting numerous additional conditions. However, this problem becomes significantly simpler when approached from the perspective of value adjustments. By reducing the magnitude of larger absolute values and increasing smaller absolute values in a bimodal distribution, all values can be shifted closer to two central points, effectively creating a double-bell-like distribution. We can demonstrate that this approach reduces the quantization loss introduced by binarization, thereby supporting the suitability of double-bell distributions for this purpose.
Theorem 2.
Given an input calibration activation $x\in\mathbb{R}^{n\times 1}$ and a weight vector $\boldsymbol{w}\in\mathbb{R}^{n\times 1}$, where $w_{i}$ is extracted from the weight matrix $\boldsymbol{W}\in\mathbb{R}^{n\times n}$ along a specific channel, we define the weight vector $\boldsymbol{w}$ as the union of two sets: - A set of several outliers with large absolute values, denoted as $U_{o}=\{o_{1}^{*},o_{2}^{*},\dots,o_{k}^{*}\}$, where $|o_{i}^{*}|\gg 0$ for $i\in\{1,\dots,k\}$; - A set of normal values with small absolute values, denoted as $U_{n}=\{n_{1},n_{2},\dots,n_{n-k}\}$, where $|n_{j}|\approx 0$ for $j\in\{1,\dots,n-k\}$. $\boldsymbol{w}=U_{o}\cup U_{n}.$ We now define a new weight vector $\boldsymbol{w}_{\text{new}}$ as follows:
  • $\boldsymbol{w}_{\text{new}}=[n_{1},n_{2},\dots,\gamma o_{1}^{*},\gamma o_{2}^{*},\dots,\gamma o_{k}^{*},\dots,n_{n-k}]$, where $\gamma\in(\frac{1}{2},1)$.
  • Alternatively, $\boldsymbol{w}_{\text{new}}=[\eta n_{1},\eta n_{2},\dots,o_{1}^{*},o_{2}^{*},\dots,o_{k}^{*},\dots,\eta n_{n-k}]$, where $\eta\in(1,2)$.
Then, the quantization error induced by $\boldsymbol{w}_{\text{new}}$, defined as $\|x\cdot\boldsymbol{w}-x\cdot\text{binarized}(\boldsymbol{w}_{\text{new}})\|$, is strictly smaller than the original quantization error $\|x\cdot\boldsymbol{w}-x\cdot\text{binarized}(\boldsymbol{w})\|$ in both cases.
Proof.
A.4.1 Scaling Up Small Values Reduces Quantization Error
Consider the scenario where the input vector is $X=[2,2,2,\dots,2]_{n}$. Assume the weights in a single channel, $W$, are given by $[\alpha_{1},\alpha_{2},...,\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $[\alpha_{i}\in U_{1}]$ is the set of values with very large absolute magnitudes, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $[\beta_{i}\in U_{2}]$ are values with magnitudes close to zero, satisfying $\sum_{i=1}^{n-k}\beta_{i}=0$. $W=U_{1}\cup U_{2}$. This structure is common in practice, as weight distributions in neural networks often exhibit a few dominant values and many small ones. As we observe, the distribution of weights along the channel dimension is mostly not symmetric around zero. Instead, it tends to be biased, with the majority leaning either towards positive or negative values. Consequently, the extreme values are predominantly either entirely positive or entirely negative, so we assume $\alpha_{i}>0$.
The product of the input $X$ and the weight vector $W$ is:
$X\cdot W^{T}=2(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k})=2(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})=2A$ (7)
since the sum of the $\beta_{i}$ is zero.
According to the quantization function, the mean value $M$ is:
$M=\frac{\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}}{n}=\frac{A}{n}$ (8)
The absolute mean value, $AbsMean$, is defined as:
$\begin{split}AbsMean&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|\beta_{1}-\frac{A}{n}|+|\beta_{2}-\frac{A}{n}|+\dots+|\beta_{n-k}-\frac{A}{n}|}{n}\end{split}$ (9)
Given that $[\beta_{1},\beta_{2},\dots,\beta_{n-1}]$ are all very small in magnitude compared to $\frac{\alpha}{n}$, we can simplify the above as:
$\begin{split}AbsMean&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-\beta_{1})+(\frac{A}{n}-\beta_{2})+\dots+(\frac{A}{n}-\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{2kA}{n})}{n}\end{split}$ (10)
Here, the sum of the $\beta_{i}$ vanishes due to their zero sum constraint.
Therefore, the dequantized value for $\alpha$ is:
$AbsMean+M=\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (11)
and for each $\beta_{i}$:
$-AbsMean+M=-\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (12)
To reduce the quantization error associated with the small-magnitude weights, we can scale them up by a factor $m>1$, while correspondingly scaling down the input. Specifically, we multiply each $\beta_{i}$ by $m$ and divide the associated input elements by $m$. The new input becomes $X_{new}=[2,2,\dots,2,\frac{2}{m},\frac{2}{m},\dots,\frac{2}{m}]_{n}$, and the new weights are $W_{new}=[\alpha_{1},\alpha_{2},...,\alpha_{k},m\beta_{1},m\beta_{2},\dots,m\beta_{n-k}]$.
The output remains unchanged:
$X_{new}\cdot W_{new}^{T}=2(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})+\frac{2}{m}\cdot m\beta_{1}+\frac{2}{m}\cdot m\beta_{2}+\dots+\frac{2}{m}\cdot m\beta_{n-k}=2A$ (13)
This invariance is crucial: the scaling operation does not affect the original computation, but it can impact the quantization error.
For the new weights, the mean value is:
$M_{new}=\frac{\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+m\beta_{1}+m\beta_{2}+\dots+m\beta_{n-k}}{n}=\frac{A+m(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}=\frac{A}{n}$ (14)
since the sum of the $\beta_{i}$ is zero.
The new absolute mean value is:
$\begin{split}AbsMean_{new}&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|m\beta_{1}-\frac{\alpha}{n}|+|m\beta_{2}-\frac{\alpha}{n}|+\dots+|m\beta_{n-k}-\frac{\alpha}{n}|}{n}\end{split}$ (15)
If $m$ is chosen such that $\frac{\alpha}{n}$ remains larger than all $m\beta_{i}$, the simplification proceeds as before:
$\begin{split}AbsMean_{new}&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-m\beta_{1})+(\frac{A}{n}-m\beta_{2})+\dots+(\frac{A}{n}-m\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-m(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{kA}{n})}{n}\end{split}$ (16)
Thus, $AbsMean_{new}$ and $M_{new}$ are identical to $AbsMean$ and $M$, and the dequantized values are unchanged. This demonstrates that scaling up the small weights does not affect the mean or absolute mean, but it can improve the quantization error, as we analyze next.
Let us now analyze the quantization error. The original output is $2\alpha$. For convenience, let $N=(n-k)(-AbsMean+M)$. The quantized output is a sum of the dequantized values for all weights.
$A>0$ Both before and after scaling, the quantization output contains the term:
$2k(AbsMean+M)=2k\left(\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}\right)$ (17)
For $n$ typically greater than 100 and $k\ll n$ it holds that:
$0<2k(AbsMean+M)<2A$ (18)
This is because the quantized value is always less than the original due to the averaging effect.
For the term $-AbsMean+M$:
$\begin{split}-AbsMean+M&=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}\\ &=-\frac{A}{n}+\frac{2kA}{n^{2}}\\ &=\frac{A}{n}\left(\frac{2k}{n}-1\right)\end{split}$ (19)
Since $n>100$ ,$A>0$ and $k\ll n$ this value is negative and its magnitude is small.
The quantization output before scaling is $2k(AbsMean+M)+2N$, and after scaling is $2k(AbsMean+M)+N$. Because $N<0$ and $2(AbsMean+M)<2A$, we have:
$2k(AbsMean+M)+2N<2k(AbsMean+M)+N<2A$ (20)
This shows that scaling up the small weights reduces the quantization error, as the quantized output moves closer to the original value.
if $\alpha_{i}<0$, the proof is similar.
In summary, scaling up the small weights (and correspondingly scaling down the input) does not change the original computation, but it systematically reduces the quantization error by making the quantized output more faithful to the original.
A.4.2 Scaling Down Large Values Reduces Quantization Error
Now, let us consider the scenario where we scale down the large-magnitude weight. Let the input $X=[1,1,1,\dots,1]_{n}$, and the weights $W=[m\alpha_{1},m\alpha_{2},...,m\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are large values, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\alpha_{i}>0$, $[\beta_{1},\dots,\beta_{n-k}]$ are small, and $\sum_{i=1}^{n-k}\beta_{i}=0$.
The output is:
$X\cdot W^{T}=m(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}=mA$ (21)
The mean value is:
$M=\frac{m\alpha_{1}+m\alpha_{2}+\dots+m\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}}{n}=\frac{mA}{n}$ (22)
The absolute mean is:
$\begin{split}AbsMean&=\frac{|m\alpha_{1}-\frac{mA}{n}|+|m\alpha_{2}-\frac{mA}{n}|+\dots+|m\alpha_{k}-\frac{mA}{n}|+|\beta_{1}-\frac{mA}{n}|+|\beta_{2}-\frac{mA}{n}|+\dots+|\beta_{n-k}-\frac{mA}{n}|}{n}\end{split}$ (23)
Since $\frac{mA}{n}$ is much larger than the $\beta_{i}$, we can simplify:
$\begin{split}AbsMean&=\frac{(m\alpha_{1}-\frac{mA}{n})+(m\alpha_{2}-\frac{mA}{n})+\dots+(m\alpha_{k}-\frac{mA}{n})+(\frac{mA}{n}-\beta_{1})+(\frac{mA}{n}-\beta_{2})+\dots+(\frac{mA}{n}-\beta_{n-k})}{n}\\ &=\frac{m(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})+(n-2k)\frac{mA}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-1})}{n}\\ &=\frac{mA+(n-2k)\frac{mA}{n}}{n}\\ &=\frac{2m(A-\frac{kA}{n})}{n}\end{split}$ (24)
The dequantized value for $m\alpha$ is:
$AbsMean+M=\frac{2m(A-\frac{kA}{n})}{n}+\frac{mA}{n}=\frac{3mA-\frac{2mkA}{n}}{n}$ (25)
and for each $\beta_{i}$:
$-AbsMean+M=-\frac{2m(A-\frac{A}{n})}{n}+\frac{mA}{n}=\frac{-mA+\frac{2mkA}{n}}{n}$ (26)
To scale down the large value $m\alpha$, we divide it by $m$ ($m>1$) and multiply the corresponding input element by $m$. The new input is $X_{new}=[m,1,1,\dots,1]_{n}$, and the new weights are $W_{new}=[\alpha_{1},\alpha_{2},\dots,\alpha_{k},\beta_{1},\beta_{2},\dots,\beta_{n-k}]$.
The output remains unchanged:
$X_{new}\cdot W_{new}^{T}=m(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}=mA$ (27)
For the new weights, the mean is:
$M_{new}=\frac{\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}}{n}=\frac{A}{n}$ (28)
The new absolute mean is:
$\begin{split}AbsMean_{new}&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|\beta_{1}-\frac{A}{n}|+|\beta_{2}-\frac{A}{n}|+\dots+|\beta_{n-1}-\frac{A}{n}|}{n}\end{split}$ (29)
With appropriate $m$, we have:
$\begin{split}AbsMean_{new}&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{\alpha}{n}-\beta_{1})+(\frac{\alpha}{n}-\beta_{2})+\dots+(\frac{\alpha}{n}-\beta_{n-k})}{n}\\ &=\frac{(\alpha_{1}+\alpha_{2}++\dots+\alpha_{k})+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-1})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{kA}{n})}{n}\end{split}$ (30)
The dequantized value for $\alpha_{i}$ is:
$AbsMean_{new}+M_{new}=\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{3A-\frac{2kA}{n}}{n}$ (31)
and for each $\beta_{i}$:
$-AbsMean_{new}+M_{new}=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{-A+\frac{2kA}{n}}{n}$ (32)
Let us now examine the quantization error. The original output is $m\alpha$. The quantization output before scaling is:
$\begin{split}&k(AbsMean+M)+(n-k)(-AbsMean+M)\\ &=k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-mA+\frac{2mkA}{n}}{n}\end{split}$ (33)
The quantization output after scaling is:
$\begin{split}&km(AbsMean_{new}+M_{new})+(n-k)(-AbsMean_{new}+M_{new})\\ &=km\frac{3A-\frac{2kA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &=k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\end{split}$ (34)
Because $A>0$, for $n>100$ ,$\alpha>0$ and $k\ll n$ it holds that:
$0<k\frac{3mA-\frac{2mkA}{n}}{n}<mA$ (35)
and
$\frac{-A+\frac{2kA}{n}}{n}=\frac{A}{n}\left(\frac{2k}{n}-1\right)<0$ (36)
Comparing the quantization outputs, we see:
$\begin{split}&k\frac{3mA-\frac{2mkA}{n}}{n}+m(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &<k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &<mA\end{split}$ (37)
If $\alpha_{i}<0$, the proof is similar.
A.5 Perplexity on C4 dataset
As shown in Tab.5, we compare the perplexity of the OPT and LLaMA families across different model sizes on C4 dataset.
[…]
Method Activation Bits OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-1-7B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
Full Precision 16 16.07 14.34 12.71 7.34 07.26 06.73 5.71
RTN 16 $9$.56 23492.89 9617.07 194607.78 115058.76 46250.21 314504.09
GPTQ 16 6364.65 6703.36 5576.82 186229.5 67954.04 19303.51 13036.32
PB-LLM 16 168.12 222.15 104.78 76.63 80.69 184.67 NAN
BiLLM 16 64.14 44.77 42.13 46.96 39.38 25.87 17.30
[…]
BiLLM 8 74.56 61.99 40.91 47.13 40.91 21.45 17.72
[…]
BiLLM 6 7348 13445.21 63.41 61.65 63.41 37.66 19.43
[…]
A.6 Visualization of distribution of activation before and after DBellQuant
In this section, we present the changes in the distribution of activation values before and after applying DBellQuant as shown in Fig.8. It is evident that the extreme values in the activation have been significantly reduced by a factor of 5 to 10; for instance, the maximum value decreases from approximately 3 to around 0.4. Previous studies have highlighted that one of the primary challenges in low-bit quantization of activations lies in the presence of large outliers, which expand the activation range and, consequently, amplify quantization errors. By applying DBellQuant, the activation range is effectively compressed from [-3, 3] to [-0.4, 0.4], dramatically alleviating the difficulty of quantization. This reduction in range establishes highly favorable conditions for further exploration of lower-bit quantization, such as 8-bit or even 6-bit implementations.
Figure 8: Visualization results of distribution of activation before and after DBellQuant across different blocks.

### v2

@@
 To address these challenges, we introduce DBellQuant, an innovative post-training quantization (PTQ) framework that achieves nearly 1-bit weight compression and 6-bit activation quantization with minimal performance degradation.
-DBellQuant uses Learnable Transformation for Dual-Bell (LTDB) algorithm, which transforms single-bell weight distributions into dual-bell forms to reduce binarization errors and applies inverse transformations to smooth activations.
+DBellQuant uses learnable transformation to map single-bell weight distribution to dual-bell distribution to reduce binarization error and smooth activations using inverse transformation.
 DBellQuant sets a new state-of-the-art by preserving superior model performance under aggressive weight and activation quantization.
-For example, on the Wikitext2 dataset, DBellQuant achieves a perplexity of 14.39 on LLaMA2-13B with 6-bit activation quantization, significantly outperforming BiLLM’s 21.35 without activation quantization, underscoring its potential in compressing LLMs for real-world applications.
+For example, on the Wikitext2 dataset, DBellQuant achieves a perplexity of 14.39 on LLaMA2-13B with nearly 1-bit weight and 6-bit activation quantization, significantly outperforming BiLLM’s 21.35 without activation quantization, underscoring its potential in compressing LLMs for real-world edge applications.
+Our code is available at: https://github.com/ZijianYY/DBellQuant.
+Machine Learning, ICML
 1 Introduction
-In recent years, the rapid advancement of large language models (LLMs) demonstrate exceptional performance in a variety of complex tasks that involve natural language understanding and generation achiam2023gpt ; dubey2024llama .
-However, these models often comprise hundreds of billions of parameters, posing significant challenges for their deployment in real-world applications because of the substantial computational and memory requirements (e.g.
+In recent years, the rapid advancement of large language models (LLMs) demonstrate exceptional performance in a variety of complex tasks that involve natural language understanding and generation (achiam2023gpt; dubey2024llama).
+However, these models often comprise hundreds of billions of parameters, posing significant challenges for their deployment in real-world edge applications because of the substantial computational and memory requirements (e.g.
 a 70B model requires around 150GB GPU memory), resulting in huge operational costs and unacceptable inference latency.
-In this context, quantization, as an effective model compression technique, has garnered significant attention.
-Due to the sparsity of information density in the weights of LLMs wang2024bitnet ; yu2024super , weight binarization has emerged as a promising quantization scheme.
-Methods such as PB-LLM shang2023pb and BiLLM huang2024billm have significantly reduced the errors introduced by binarization by applying finer-grained processing to critical weights.
-Despite its theoretical advantages, the application of quantization in compressing LLMs continues to face numerous challenges egashira2024exploiting .
-For instance, systematic outliers in activation values can lead to substantial quantization errors, negatively impacting model accuracy sun2024massive .
-Although recent studies xiao2023smoothquant ; shao2023omniquant have proposed mitigating such errors by redistributing the scaling factors between weights and activation values.
-Other research directions ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm focus on leveraging Hadamard transformations to effectively eliminate outlier features.
-However, no binarization methods have taken care of simultaneous smoothing activation outliers.
-Consequently, the demand of higher bit-width activations during computation leads to significant computational overhead.
-BitNet a4.8 wang2024bitnet addresses this issue by employing quantization-aware training (QAT) to achieve a 1-bit LLM with 4-bit activations.
-Nevertheless, the QAT approach requires extensive computational resources and prolonged training time.
-Therefore, developing post-training quantization (PTQ) methods for binary compression that minimize performance loss while simultaneously smoothing activation and thus reducing activation bit-width remains a critical and challenging task.
-Figure 1: Performance on Wikitext2 dataset. DBellQunat outperforms weight-only quantization method under 8-bit activation setting.
+Figure 1: Performance on Wikitext2 dataset. DBellQuant outperforms weight-only quantization method under 8-bit activation setting.
+Figure 2: (a) Before applying DBellQuant, activations exhibit significant outliers, making quantization challenging, while the single-bell-shaped weight distribution hinders binarization.
+(b) After applying DBellQuant, activations are smoothed with substantially fewer outliers, facilitating easier quantization.
+Weight distribution is transformed to dual-bell form, which is more conducive to binarization.
+Quantization is a compelling route to LLM compression, and weight binarization is especially attractive for sharply reducing model size (wang2024bitnet; yu2024super).
+Recent PTQ methods– e.g., PB-LLM and BiLLM– mitigate binarization error with finer-grained treatment of salient weights (shang2023pb; huang2024billm).
+However, activation outliers remain a major obstacle to simultaneously quantizing activations (sun2024massive), undercutting memory and latency gains.
+Scale-redistribution techniques (xiao2023smoothquant; shao2023omniquant) and Hadamard-based transforms (ashkboos2024quarot; liu2024spinquant; sun2025flatquantflatnessmattersllm; lin2024duquantdistributingoutliersdual) help at higher weight precisions, but no PTQ weight binarization method concurrently achieves effective activation quantization.
+This is because pushing weights to 1-bit increases sensitivity to activation outliers, while existing fixes shift the difficulty back to the weights or require higher-bit activations for accuracy.
+Quantization-aware training (QAT) approaches such as BitNet can reach 1-bit weights with low-bit activations (wang2024bitnet), but at the cost of heavy retraining.
+Thus, there is a need for post-training weight binarization that preserves accuracy while also enabling activation quantization.
 To this end, we begin by revisiting the distribution characteristics of activations and weights in LLMs (Fig.
@@
 2(b)) can effectively reduce binarization errors.
+In addition, activation distribution suffer from outliers and thus large quantization errors, which demands a narrower distribution.
 This motivates a key question: Is it possible to transform weights into a dual-bell shape distribution while simultaneously addressing activation outliers to facilitate both activation quantization and weight binarization?
@@
 DBellQuant enables activation quantization while achieving near 1-bit weight compression with minimal accuracy loss.
-At the core of DBellQuant is a learnable transformation matrix that maps the weight distribution into a dual-bell form, while its inverse is applied to the input activations to smooth them.
-We begin by analyzing the characteristics of weight distributions suitable to binarization and theoretically derive the feasibility of applying a dual-bell transformation.
-Based on this, we then develop an efficient and lightweight algorithm, Learnable Transformation for Dual-Bell (LTDB).
-LTDB initializes the transformation matrix using an activation-aware strategy and optimizes it via a custom objective function that encourages the weights to cluster around two centers.
-This drives the formation of a symmetric dual-peak distribution, and an early stopping strategy is employed to ensure stable and efficient optimization.
-Through this equivalence-preserving transformation, weights originally exhibiting unimodal distributions– challenging for binarization– are mapped into near-symmetric dual-bell distributions (Fig.
+The core mechanism is a learnable, equivalence-preserving transformation that maps each layer’s weight distribution toward a dual-bell shape; its inverse is applied to the input activations to keep the computation unchanged.
+We analyze the feasibility of dual-bell transformations for binarization and propose a lightweight algorithm, Learnable Transformation for Dual-Bell (LTDB), which initializes the transform and optimizes it with a custom objective that drives weights to cluster around two modes, with early stopping to ensure stability.
+By doing so, weights originally exhibiting unimodal distributions– challenging for binarization– are mapped into near-symmetric dual-bell distributions (Fig.
 2(b)), significantly reducing binarization error.
-Simultaneously, the inverse of the learned transformation is applied to the input activations, effectively scaling down outliers and smoothing the distribution.
-This makes the activations more amenable to quantization without altering the model’s output (Fig.
+Moreover, as discussed in Sec.
+3.4, experiments show that applying the inverse transform to inputs consistently contracts activation ranges and suppresses outliers, making activations more amenable to quantization (Fig.
 2(b)).
 Experimental results demonstrate that the equivalent transformation strategy of DBellQuant significantly reduces the loss caused by weight binarization.
-For the first time under PTQ conditions, it achieves near 1-bit weight compression while simultaneously compressing activations to 6 bits.Across various LLMs and evaluation metrics, DBellQuant consistently achieves state-of-the-art results.
-For instance, on the Wikitext2 benchmark, we achieves a perplexity of 14.39 on LLaMA2-13B using only 6-bit activations (Fig.
-1), significantly surpassing the performance of BiLLM, a method that only quantizes weights, which achieves a perplexity of 21.35.
-Figure 2: (a) Before applying DBellQuant, activations exhibit significant outliers, making quantization challenging, while the single-bell-shaped weight distribution hinders binarization.
-(b) After applying DBellQuant, activations are smoothed with substantially fewer outliers, facilitating easier quantization.
-Weight distribution is transformed to dual-bell form, which is more conducive to binarization.
+For the first time under PTQ conditions, it achieves near 1-bit weight compression while simultaneously compressing activations to 6 bits or even 4bits.
+Across various LLMs and evaluation metrics, DBellQuant consistently achieves state-of-the-art results.
+For instance, on the Wikitext2 benchmark, we achieve a perplexity of 21.69 on LLaMA-2-7B using only 6-bit activations, even with 4-bit activations, we achieve a perplexity of 23.04, significantly surpassing the performance of BiLLM, which achieves a perplexity of 21.35 .
 2 Related Work
@@
 LLMs quantization have introduced a variety of innovative techniques to enhance efficiency while maintaining accuracy.
-Works like GPTQ frantar2022gptq and OBQ frantar2023optimalbraincompressionframework minimizes reconstruction error by adjusting the remaining unquantized parameters in the block to compensate for the accuracy loss caused by quantization.
-LLM.int8()dettmers2022gpt3 and ZeroQuantyao2022zeroquant improve quantization accuracy by introducing additional grouping labels for customized quantization blocks.
-Other works like SmoothQuantxiao2023smoothquant and OmniQuant shao2023omniquant addresses activation outliers by redistributing scaling factors between weights and activations, migrating the quantization difficulty from activation to weights.
-Additionally, recent approaches leverage Hadamard transformations to suppress activation outliers ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm , while incoherence processing has been proposed for effective low-bit quantization chee2024quip2bitquantizationlarge ; tseng2024quipbetterllmquantization .
-Collectively, these advancements demonstrate that quantization techniques can be successfully scaled to multi-billion-parameter models, achieving substantial reductions in memory consumption and inference latency without compromising model performance.
+Works like GPTQ (frantar2022gptq) and OBQ (frantar2023optimalbraincompressionframework) minimizes reconstruction error by adjusting the remaining unquantized parameters in the block to compensate for the accuracy loss caused by quantization.
+LLM.int8()(dettmers2022gpt3) and ZeroQuant(yao2022zeroquant) improve quantization accuracy by introducing additional grouping labels for customized quantization blocks.
+Other works like SmoothQuant(xiao2023smoothquant) and OmniQuant (shao2023omniquant) addresses activation outliers by redistributing scaling factors between weights and activations, migrating the quantization difficulty from activation to weights.
+Additionally, recent approaches leverage Hadamard transformations to suppress activation outliers (ashkboos2024quarot; liu2024spinquant; sun2025flatquantflatnessmattersllm), while incoherence processing and non-linear transformation have been proposed for effective low-bit quantization (chee2024quip2bitquantizationlarge; tseng2024quipbetterllmquantization; zhang2024magrweightmagnitudereduction).
+Collectively, these advancements demonstrate that quantization techniques can be scaled to multi-billion-parameter models, achieving substantial reductions in memory consumption and inference latency without compromising model performance.
 Binary Quantization
@@
 However, applying binary quantization to LLMs presents substantial challenges due to their sensitivity to precision loss, particularly in attention mechanisms and large embedding layers.
-BinaryBERT bai2020binarybert explored binary quantization for BERT, proposing selective preservation of critical weights in higher precision to mitigate performance degradation.
-In another direction, PB-LLM shang2023pb introduced a partially-binarized approach for LLMs, retaining a small fraction of salient weights in higher precision while binarizing the rest, enabling extreme low-bit quantization without sacrificing linguistic reasoning capabilities.
-Recent advancements include structural binarization techniques that leverage novel sparsity forms and standardized importance metrics to selectively binarize and sparsify LLM weights dong2024stbllmbreaking1bitbarrier , as well as strategies like alternating refined binarization and column-group bitmap methods to effectively reduce quantization error and address column deviations li2024arbllmalternatingrefinedbinarizations .
+BinaryBERT (bai2020binarybert) explored binary quantization for BERT, proposing selective preservation of critical weights in higher precision to mitigate performance degradation.
+In another direction, PB-LLM (shang2023pb) introduced a partially-binarized approach for LLMs, retaining a small fraction of salient weights in higher precision while binarizing the rest, enabling extreme low-bit quantization without sacrificing linguistic reasoning capabilities.
+Recent advancements include structural binarization techniques that leverage novel sparsity forms and standardized importance metrics to selectively binarize and sparsify LLM weights (dong2024stbllmbreaking1bitbarrier), as well as strategies like alternating refined binarization and column-group bitmap methods to effectively reduce quantization error and address column deviations (li2024arbllmalternatingrefinedbinarizations).
 These innovations collectively advance the feasibility of binary quantization for LLMs, pushing the boundaries of efficiency without compromising performance.
@@
 (d)The weight distribution after transformation will be double-bell.
-3 Method
+3 Methodology
 We begin by exploring the process of binarization, analyzing and theoretically proving the weight distributions suitable for binarization in Sec.
@@
 By utilizing the sign function, binarization can convert weights in LLMs into binary values. The per-channel binarization and de-binarization process is as follows:
-$\small\beta=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{W_{i,j}},\quad\boldsymbol{\widetilde{W}}=\text{Sign}(\boldsymbol{W}-\beta),\quad\alpha=\frac{1}{n}\sum_{i=1}^{n}|\boldsymbol{W_{i,j}}-\beta_{j}|$ (1)
-$\text{Sign}(\boldsymbol{W_{i,j}})=\begin{cases}+1,&\text{if }\boldsymbol{W_{i,j}}>0,\\ -1,&\text{if }\boldsymbol{W_{i,j}}\leq 0,\end{cases},\quad\boldsymbol{W_{deq}}=\boldsymbol{\widetilde{W}}\cdot\alpha+\beta$ (2)
+$\tiny\beta=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{W_{i,j}},\quad\boldsymbol{\widetilde{W}}=\text{Sign}(\boldsymbol{W}-\beta),\quad\alpha=\frac{1}{n}\sum_{i=1}^{n}|\boldsymbol{W_{i,j}}-\beta_{j}|$ (1)
+$\tiny\text{Sign}(\boldsymbol{W_{i,j}})=\begin{cases}+1,&\text{if }\boldsymbol{W_{i,j}}>0,\\ -1,&\text{if }\boldsymbol{W_{i,j}}\leq 0,\end{cases},\quad\boldsymbol{W_{deq}}=\boldsymbol{\widetilde{W}}\cdot\alpha+\beta$ (2)
 where $\beta$ is the shifting factor and $\alpha$ is the scaling factor for binarization.
-Previous studies huang2024slim have shown that neural network weights exhibit structured distributions along the channel dimension, with certain channels being more salient.
-The overall weight distribution typically follows a quasi-Gaussian pattern dettmers2022llmint88bitmatrixmultiplication , as does the channel-wise distribution (Fig.
+Previous studies (huang2024slim) have shown that neural network weights exhibit structured distributions along the channel dimension, with certain channels being more salient.
+The overall weight distribution typically follows a quasi-Gaussian pattern (dettmers2022llmint88bitmatrixmultiplication), as does the channel-wise distribution (Fig.
 6).
@@
 However, LLM weight distributions typically exhibit single-bell characteristics, and conventional PTQ methods fail to effectively transform them for binarization.
-While QAT can reshape weight distributions into a dual-bell form through its learning objectives wang2023bitnetscaling1bittransformers , it requires substantial computational resources and prolonged training.
+While QAT can reshape weight distributions into a dual-bell form through its learning objectives (wang2023bitnetscaling1bittransformers), it requires substantial computational resources and prolonged training.
 To address this, we propose a more efficient PTQ method that rapidly converts single-bell distributions into dual-bell ones, optimizing binary quantization without the need for resource-intensive retraining.
 3.2 Learnable Transformation for Dual-Bell Quantization
-Learnable Transformation with Auxilary Matrix
+Learnable Transformation with Auxiliary Matrix
 As mentioned before, double-bell distributions are advantageous for binarization.
@@
 $\boldsymbol{w}_{i}^{\prime}\sim\pi\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(\mu_{2},\sigma_{2}^{2}),$
-where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are parameters of the doubel-bell distribution. More detailed proof is shown in Appendix. A.3.
+where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are parameters of the double-bell distribution. More detailed proof is shown in Appendix. A.19.
 Theorem 1 demonstrates that transforming weight distributions from single-bell to double-bell can be achieved by introducing an auxiliary matrix $T$.
@@
 All values can be shifted closer to two central points through these two processes, and it will reduce the quantization error shown in Appendix.
-A.4.
+A.20.
 Regarding activation quantization, the initialization explicitly accounts for outliers in the activation matrix, making it inherently activation-aware.
@@
 Our learning objective is to encourage all weight values to move closer to the two mean centers calculated by Eq.
-2, ultimately forming a doubel-bell distribution with these two values as its peaks.
+2, ultimately forming a double-bell distribution with these two values as its peaks.
 During the binarization process, we denote these two points as $m_{1}$ and $m_{2}$ respectively.
@@
 As discussed before, DTMD alone is not enough to train a better weights distribution for binarization. So we introduce a new loss function as follows:
-$\small\mathcal{L}_{\text{DTNP}}=\frac{\lambda_{DTNP}}{n}\sum_{i=1}^{n}\begin{cases}\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|}{\left|\text{m}_{1,i}\right|},&\text{if }\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|<\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|\\[10.0pt] \frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|}{\left|\text{m}_{2,i}\right|},&\text{otherwise.}\end{cases}$ (6)
+$\tiny\mathcal{L}_{\text{DTNP}}=\frac{\lambda_{DTNP}}{n}\sum_{i=1}^{n}\begin{cases}\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|}{\left|\text{m}_{1,i}\right|},&\text{if }\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|<\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|\\[10.0pt] \frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|}{\left|\text{m}_{2,i}\right|},&\text{otherwise.}\end{cases}$ (6)
 Figure 4: Dual-Target Minimum Deviation Loss value over iterations across different layers.
@@
 3.4 Impact of the Inverse of Learnable Transformation Matrix on Activation Smoothing
-Figure 5: Visualization of distribution of values in $T$ and $T^{-1}$ of Llama2-7B.
-Through the use of DTNP for training, we understand that the transformation matrix $T$ drives all values in the weights closer to the two mean centers calculated by Eq.
+Driven by the DTNP objective, the transformation matrix $T$ encourages weight values to cluster around the two binarization centers defined in Eq.
 2.
-Prior research has shown that the distribution of weights tends to approximate a quasi-Gaussian distribution, with the majority of values being extremely small and close to zero, while only a very small fraction exhibit relatively large absolute values dettmers2022llmint88bitmatrixmultiplication .
-Theoretically, this implies that $T$ will contain many values greater than 1 to amplify the numerous near-zero absolute values, bringing them closer to the mean centers.
-At the same time, very few values of $T$ will be less than 1 to reduce the relatively rare large absolute values, aligning them similarly with the mean centers.
-In fact, when we visualize $T$ after training as shown in Fig.
-5, we observe that less than 5% of its values are below 1.
-Consequently, the corresponding $T^{-1}$, which is multiplied with activations, has over 95% of its values below 1.
-This significantly reduces the quantization range of activations and suppresses the magnitude of outlier values within the activations.
-Visualization results can be seen in Appendix.
-A.6.
-As a result, this approach is particularly effective in further facilitating activation quantization.
+Given that LLM weights typically follow a quasi-Gaussian distribution concentrated near zero (dettmers2022llmint88bitmatrixmultiplication), $T$ tends to learn values greater than 1 to amplify these small magnitudes towards the binarization targets.
+Crucially, this mechanism inherently benefits activation quantization.
+Our activation-aware initialization (Eq.
+4) establishes a positive correlation between the activation magnitude $\max(|X_{j}|)$ and the transformation scale $T_{j}$.
+Consequently, channels exhibiting significant activation outliers are assigned larger $T_{j}$ values.
+When the inverse transformation $T^{-1}$ is applied to the input $X$, these outlier-heavy channels undergo stronger suppression (multiplication by a smaller $1/T_{j}$ factor) compared to normal channels.
+Empirically, as visualized in Fig.
+5, we observe that less than 5% of values in $T$ are below 1, implying that $T^{-1}$ acts as a global compressor with channel-specific intensity.
+This effectively narrows the dynamic range of activations and mitigates the kurtosis caused by outliers.
+We provide a quantitative validation of this smoothing effect in Appendix A.17, where metrics including Z-score and Relative Deviation Error demonstrate a substantial reduction in distribution variance and outlier impact, making the transformed activations significantly more amenable to low-bit quantization.
 3.5 Algorithm
 The process of Learnable Transformation for Dual-Bell Transformation Algorithm is shown in Algorithm.
-1, which adjusts the full-precision weight matrix $W$ using an activation-aware initialization transformation matrix $T$ over $N$ epochs.
+1, which can be seen in Appendix.
+A.18 and can adjust the full-precision weight matrix $W$ using an activation-aware initialization transformation matrix $T$ over $N$ epochs.
 The algorithm iteratively minimizes dual-target loss functions to guide $W$ toward a dual-bell distribution while employing an early stopping mechanism based on the DTMD.
-Algorithm 1 Learnable Transformation for Dual-Bell Transformation (LTDB)
-1:function LTDB($\boldsymbol{W},\mathbf{T},N$)
-2:     Input: $\boldsymbol{W}\in\mathbb{R}^{n\times m}$ a full-precision weight matrix.
-3:     Input: $\mathbf{T}\in\mathbb{R}^{1\times m}$ an activation-aware initialization transformation matrix.
-4:     Input: $N$ the total number of epochs.
-5:     Output: $\widetilde{\boldsymbol{W}}\in\mathbb{R}^{n\times m}$ the transformed weight matrix.
-6:     for $\text{iter}=1,2,\dots,N$ do
-7:         $\widetilde{\boldsymbol{W}}\leftarrow\mathbf{T}\odot\boldsymbol{W}$triangleright Perform element-wise multiplication
-8:         $\mathcal{L}_{\text{DTMD}},\mathcal{L}_{\text{DTNP}}\leftarrow\text{LossFunc}(\widetilde{\boldsymbol{W}})$triangleright Compute the two dual-target losses for $\widetilde{\boldsymbol{W}}$
-9:         $\mathcal{L}_{\text{DTNP}}.backward()$triangleright Use Dual-Target Normalized Proportional Loss for training
-10:         if $\mathcal{L}_{\text{DTMD}}>\mathcal{L}_{\text{DTMD-Minimum}}$ then
-11:              break $\triangleright$ Stop training based on Dual-Target Minimum Deviation Loss
-12:         end if
-13:         $\mathcal{L}_{\text{DTMD-Minimum}}\leftarrow\mathcal{L}_{\text{DTMD}}$
-14:     end for
-15:     return $\widetilde{\boldsymbol{W}}$
-16:end function
 4 Experiments
 4.1 Settings
-All experimental procedures were executed utilizing the PyTorch paszke2019huggingface framework in conjunction with the Huggingface library paszke2019huggingface .
+All experimental procedures were executed utilizing the PyTorch (paszke2019huggingface) framework in conjunction with the Huggingface library (paszke2019huggingface).
 Models with parameters smaller than 8B are running on a single NVIDIA A30 GPU equipped with 24GB of memory, others are running on a single NVIDIA A100 GPU equipped with 80GB of memory.
 Consistent with methodologies outlined by Frantar et al.
-frantar2023gptq and Huang et al.
-huang2024billm , a calibration dataset comprising 128 samples sourced from the C4 collection raffel2020exploring was employed.
+(frantar2023gptq) and Huang et al.
+(huang2024billm), a calibration dataset comprising 128 samples sourced from the C4 collection (raffel2020exploring) was employed.
+Figure 5: Visualization of distribution of values in $T$ and $T^{-1}$ of Llama2-7B.
 Models and Datasets
-Comprehensive evaluations were carried out across several large language model families, including LLaMA, LLaMA-2, and LLaMA-3 touvron2023llama and the OPT series zhang2022opt .
-The efficacy of the developed DBellQuant was assessed by calculating the perplexity of the models’ generated text on standard benchmarks: WikiText2 merity2017pointer , and a subset of the C4 data raffel2020exploring .
-Furthermore, the models’ performance was evaluated based on accuracy across seven zero-shot question-answering tasks: ARC-c clark2018think , ARC-e clark2018think , Hellaswag zellers2019hellaswag , PIQA bisk2020piqa , and Winogrande sakaguchi2020winogrande .
+Comprehensive evaluations were carried out across several large language model families, including LLaMA, LLaMA-2, and LLaMA-3 (touvron2023llama), OPT series (zhang2022opt), Qwen series (qwen2025qwen25technicalreport) and Deepseek Model (Guo_2025).
+The efficacy of the developed DBellQuant was assessed by calculating the perplexity of the models’ generated text on standard benchmarks: WikiText2 (merity2017pointer), and a subset of the C4 data (raffel2020exploring).
+Furthermore, the models’ performance was evaluated based on accuracy across five zero-shot question-answering tasks: ARC-c (clark2018think), ARC-e (clark2018think), Hellaswag (zellers2019hellaswag), PIQA (bisk2020piqa), and Winogrande (sakaguchi2020winogrande).
 Comparison Methods
-The primary benchmark for comparison for our DBellQuant approach is BiLLM huang2024billm , which represents the current baseline PTQ technique for binary large language models.
-Additionally, we include other contemporary PTQ algorithms in our comparison, namely Round-to-Nearest (RTN), GPTQ frantar2023gptq , PB-LLM shang2023pb and the current state-of-the-art PTQ technique ARB-LLM li2024arbllmalternatingrefinedbinarizations .
+The primary benchmark for comparison for our DBellQuant approach is BiLLM (huang2024billm), which represents the current baseline PTQ technique for binary large language models.
+Additionally, we include other contemporary PTQ algorithms in our comparison, namely Round-to-Nearest (RTN), GPTQ (frantar2023gptq), PB-LLM (shang2023pb) and the current state-of-the-art PTQ technique ARB-LLM (li2024arbllmalternatingrefinedbinarizations).
+Bit-width LTDB Algorithm WikiText2 $\downarrow$ C4 $\downarrow$
+16 24.41 27.69
+16 17.91 21.83
+8 27.39 30.91
+8 18.65 23.80
+6 32.74 42.48
+6 21.69 30.24
+4 34.62 45.93
+4 23.04 29.90
+Table 1: Performance w/o LTDB Algorithm.
+Method MathQA LogiQA2
+fp16 44.96 $\pm$ 0.91 29.58 $\pm$ 1.15
+GPTQ 2bit 20.51 $\pm$ 0.82 14.65 $\pm$ 1.53
+BiLLM 1.1bit 24.99 $\pm$ 0.79 20.42 $\pm$ 1.87
+DBellQuant 1.1bit 27.65 $\pm$ 0.75 22.78 $\pm$ 1.48
+Table 2: DeepSeek-R1-Distill-Qwen-7B results.
+Method TextVQA ChartQA MME-Per
+fp16 84.9 87.3 1695
+GPTQ 2bit 0 0 319
+BiLLM 1.1bit 26.3 3.7 638
+DBellQuant 1.1bit 28.4 5.2 742
+Table 3: Qwen-2.5-VL-7B results.
 […]
@@
 […]
-[…]
+DBellQuant 16 43.42 31.47 18.89 15.34 17.91 12.79 6.84
 BiLLM 8 88.95 68.60 166.46 40.13 33.23 22.55 14.72
-[…]
+DBellQuant 8 44.98 30.39 18.88 14.74 18.65 13.11 6.88
 BiLLM 6 9537 18405.85 28123.58 71.65 42.41 30.20 18.65
-[…]
+DBellQuant 6 61.50 47.33 21.12 16.66 21.69 14.39 7.56
 4.2 Main Results
 […]
+Method wikitext2 C4
+BiLLM 38.42 40.81
+DBellQuant 27.05 31.02
+Table 5: Qwen-2-7B results.
+Method wikitext2 C4
+BiLLM 41.74 53.07
+DBellQuant 26.78 30.76
+Table 6: Qwen-2.5-7B results.
+Model 4096×4096 4096×11008 11008×4096
+FP16 0.79463 1.73942 1.82653
+BiLLM 0.36842 0.38744 0.43906
+ARB-LLMX 0.33180 0.35539 0.36792
+DBellQuant 0.27694 0.30085 0.31860
+Table 7: Actual inference speed comparison on various linear layers in LLaMA-2-7B.
 Table 8: Accuracy of PIQA, ARC-e, ARC-c, HellaSwag, Winogrande and average accuracy of all datasets with BiLLM and our methods on OPT and LLaMA family.
+Values marked with * indicate usage of NVFP4 format, while other values are represented in integer format.
 Model Method Activation Bits PIQA ARC-e Arc-c HellaSwag Winogrande Avg.
@@
 BiLLM 16 59.63 36.83 17.06 30.14 51.30 38.99
+[…]
 DBellQuant 16 70.29 55.76 24.72 37.81 58.71 49.46
-[…]
+OPT-6.7B BiLLM 8 54.30 31.02 20.05 26.66 50.90 36.59
 DBellQuant 8 69.10 54.63 24.91 38.19 57.38 48.84
-[…]
+BiLLM 6 53.43 20.08 20.22 25.80 47.67 33.44
 DBellQuant 6 68.12 52.44 23.38 37.04 57.30 47.65
+DBellQuant 4* 70.24 55.72 23.37 38.01 53.43 48.81
 - 16 78.40 67.34 38.14 56.45 67.01 61.46
 BiLLM 16 61.92 38.93 21.58 32.78 53.67 41.77
+[…]
 DBellQuant 16 67.74 49.37 24.23 39.55 58.80 47.94
-[…]
+LLaMA-1-7B BiLLM 8 61.86 37.88 21.76 32.09 51.62 41.04
 DBellQuant 8 67.41 47.31 26.19 38.89 58.80 47.72
-[…]
+BiLLM 6 57.45 31.06 20.65 29.78 53.04 38.40
 DBellQuant 6 65.29 46.46 25.77 37.28 54.85 45.94
+DBellQuant 4* 62.91 40.22 23.24 32.76 52.48 42.32
 - 16 78.40 69.28 40.02 56.69 67.25 62.32
 BiLLM 16 60.34 36.87 21.59 30.24 51.62 40.13
+[…]
 DBellQuant 16 63.98 42.85 23.89 34.82 56.27 44.36
-[…]
+LLaMA-2-7B BiLLM 8 59.74 36.95 21.42 30.96 53.75 40.56
 DBellQuant 8 62.56 42.42 23.03 34.44 54.38 43.37
-[…]
-DBellQuant 6 61.10 37.5 22.18 31.94 53.85 41.32
+BiLLM 6 56.81 28.32 20.05 29.33 52.17 37.33
+DBellQuant 6 61.10 37.50 22.18 31.94 53.85 41.32
+DBellQuant 4* 60.82 38.35 22.69 32.76 53.35 41.59
 - 16 79.65 80.09 50.51 60.18 72.77 68.64
 BiLLM 16 57.51 33.75 18.52 31.63 53.12 38.90
-DBellQuant 16 62.35 44.11 19.97 33.19 55.96 43.12
 […]
+DBellQuant 16 64.15 44.11 22.02 34.73 56.36 44.27
+LLaMA-3-8B BiLLM 8 60.55 37.96 18.34 32.60 51.78 40.24
 DBellQuant 8 61.70 40.95 18.40 32.95 55.09 41.82
+BiLLM 6 56.09 33.92 16.72 31.75 51.85 38.06
+DBellQuant 6 58.00 37.63 18.77 31.83 51.92 39.64
+DBellQuant 4* 60.07 40.40 19.89 31.92 53.43 41.14
+4.3 Ablation Experiments
+Effectiveness of LTDB algorithm
+To validate the effectiveness of our advanced LTDB algorithm, we compare them with the vanilla initialization without LTDB.
+We can observe that under different bitwidths of activation quantization, the models utilizing the proposed LTDB method consistently outperform those relying on vanilla initialization as shown in Tab.
+1.
+This demonstrates that the proposed LTDB method can effectively reduce quantization errors.
+Optimization Objective
+In DTNP, we adopted L1 loss as the objective function, as it consistently outperformed L2 loss in our experiments.
+Additionally, we introduced the hyperparameter $\epsilon$ during the activation-aware initialization process and found that all tested values surpassed prior algorithms, with $\epsilon=0.85$ achieving the best performance for LLaMA-2-7B as shown in Tab.
+21 and Tab.
+16.
+Detailed parameter settings are provided in the Appendix.
+A.6.
+Impact of Block Size
+We evaluated the impact of block size on DBellQuant’s quantization performance using block sizes ranging from 64 to 256 columns, as shown in Tab.
+19.
+Similar to other PTQ methods, smaller block sizes achieve lower perplexity but increase quantization diversity and weighting overhead.
+In previous experiments, we used a block size of 128, as it balances bit-width and quantization performance effectively.
+Notably, our method consistently outperforms baseline approaches across all block sizes, highlighting its robustness and generalizability.
+Activation Bit-width Comparisons
+We compared model performance across different activation bit-widths, as shown in Fig.
+7.
+When activations are quantized to 8 bits, perplexity remains nearly unchanged or even improves.
+Even with lower-bit quantization, such as 6-bit, the performance remains largely intact, with minimal perplexity increases.
+Notably, we also conducted experiments with NVFP4 format 4-bit activation quantization (chen2025int) in Tab.
+8 and Tab.
+9, we can see the model can still retain a significant level of performance.
+In most cases, it even outperforms 16-bit BiLLM.
+These results highlight the potential of our method in enabling efficient and reliable model performance under lower-bit settings.
+Implementation on Various Advanced Models
+To demonstrate the adaptability of DBellQuant beyond standard benchmarks, we extended our evaluation to the Qwen family, reasoning, and multi-modal models.
+As shown in Tab.5 and Tab.6, DBellQuant consistently outperforms BiLLM on both Qwen-2-7B and Qwen-2.5-7B.
+Furthermore, we evaluated the reasoning model DeepSeek-R1-Distill-Qwen-7B (Tab.2) and the multi-modal Qwen-2.5-VL-7B (Tab.3).
+Despite the significant challenge of near 1-bit quantization causing noticeable performance drops, our method still outperforms BiLLM across MathQA, LogiQA2, and visual benchmarks, highlighting its robustness in ultra-low-bit scenarios.
+4.4 Time and Memory Analysis
+Time Comparison
 […]
-DBellQuant 6 58.00 37.63 18.77 31.83 51.92 39.64
-4.3 Ablation Experiments
-Optimization Objective
-In the definition of DTNP, we employed the L1 loss as described in Eq.
-6.
-To evaluate the effectiveness of the L1 loss, we tested by replacing it with L2 loss and do the same experiments.
-We observed that the overall performance was inferior to that achieved by L1 loss.
-Therefore, we selected L1 loss as the loss function for DTNP.
-Additionally, during the activation-aware initialization process, the parameter $\epsilon$ is introduced as a hyperparameter.
-We conducted experiments with different values of $\epsilon$, and the results demonstrated that all tested values of $\epsilon$ consistently outperformed prior algorithms.
-For LLaMA-2-7B, setting $\epsilon$ to 0.85 yielded the best performance.
-Results are shown in Tab.
-4.
-Impact of Block Size
-We investigated the impact of block size on the quantization performance of DBellQuant by experimenting with block sizes ranging from 64 to 256 columns.
-The results are shown in Tab.
-3.
-Consistent with other PTQ methods, smaller block sizes yield lower perplexity, but it will increase the diversity of quantization results also increase the weighting overhead.
-In all experiments shown before we set the block size to 128 because we think it can better balance the bit-width and quantization effect.
-Furthermore, our method consistently outperforms the current baseline approaches across various block sizes, demonstrating its robustness and generalizability under different configurations.
+Memory Comparison
+We include a detailed comparison of bitwidth and model size across different settings.
+Our method reduces the model size to approximately 1/7 to 1/6 of the original FP16 model, with only a minimal increase in perplexity.
+Compared to other ultra-low-bit PTQ methods such as BiLLM and PB-LLM, our approach achieves comparable bitwidth and model size, while yielding significantly lower perplexity as shown in Tab.
+23.
+More results can be seen in Appendix.
+A.12.
+5 Conclusion
+We propose DBellQuant, an efficient PTQ method enabling simultaneous weight binarization and activation quantization for LLMs.
+By leveraging weight distributions suited for binarization, we design the Learnable Transformation for Dual-Bell algorithm.
+This includes two customized loss functions and an early stopping mechanism to achieve the dual-bell transformation.
+This transformation enhances activation quantization at the same time, enabling near 1-bit weight compression and 6-bit or even 4-bit activation quantization with minimal performance loss—achieving this milestone for the first time in the PTQ framework.
+Experiments on open-source LLMs across different families and sizes demonstrate that DBellQuant significantly advances the performance of previous SOTA binary PTQ methods.
+Impact Statement
+This paper enhances inference-time and memory efficiency for LLMs by introducing DBellQuant, a novel post-training quantization framework.
+DBellQuant significantly reduces computational and memory demands through nearly 1-bit weight compression and 6-bit or even 4-bit activation quantization while maintaining model performance.
+Its contributions improve the practicality and accessibility of deploying LLMs in resource-constrained environments, such as edge devices.
+The societal and ethical implications align with standard advancements in ML efficiency, promoting sustainability and usability without introducing new model capabilities, data sources, or application domains.
+As such, it raises no additional ethical concerns beyond those already associated with LLMs and presents no new broader impacts.
+Appendix A Appendix
+A.1 Use of Large Language Models
+In preparing this manuscript, we exclusively used large language models—GPT-5 (openai2025gpt5)—to refine grammar, flow, and tone at the sentence and paragraph levels.
+These tools were not used to generate ideas, design experiments, or draw conclusions.
+All technical content, methodologies, and interpretations were independently authored, rigorously verified, and approved by the authors.
+To prevent factual inaccuracies or citation errors, every model-edited sentence was reviewed by the authors, and all references were meticulously cross-checked with their original sources.
+The authors take full responsibility for the accuracy and integrity of this work.
+A.2 Limitations
+Our current work supports quantizing activations to 6 bits and can handle certain tasks with 4-bit activation quantization while maintaining competitive performance.
+However, for more challenging tasks, the model may still encounter performance degradation, indicating the need for additional techniques to address these issues.
+Recent studies have not only demonstrated the feasibility of 4-bit activation quantization but also explored methods to mitigate its limitations.
+Inspired by these advancements, we aim to further optimize our models by combining low-bit quantization with other techniques.
+In addition, sparse attention has recently gained significant attention, with numerous studies highlighting its potential to accelerate inference (wang2025viso; li2025efficient).
+A key objective of our work is to explore the integration of sparse attention with lower-bit quantization techniques, enabling faster computation while maintaining model performance.
+This combination could open new possibilities for efficient and practical deployment in real-world scenarios.
+A.3 Broader Impacts
+Our work demonstrates the feasibility of simultaneously binarizing weights and quantizing activations to 6 bits in LLMs while maintaining competitive performance.
+This approach significantly reduces the computational and memory overhead associated with LLM deployment, making them more accessible for resource-constrained environments.
+By enabling efficient inference (yang2024conditional), our method contributes to the democratization of advanced AI technologies, reducing the environmental impact of large-scale model deployment.
+Furthermore, it opens new avenues for research in ultra-low-bit quantization, fostering innovation in model efficiency and scalability.
+A.4 Conceptual Comparison with Prior PTQ Methods
+In this section, we clarify the conceptual differences between DBellQuant and several closely related post-training quantization methods, such as SmoothQuant (xiao2023smoothquant), OmniQuant (shao2023omniquant), and QuaRot (ashkboos2024quarot).
+While all these methods can be interpreted as applying some form of pre/post-transform to weights and/or activations, their design goals, operating bit-widths, and transformation mechanisms are fundamentally different.
+Compared to these prior approaches, DBellQuant is not merely another form of scale redistribution:
+  • Distribution-shaping for binary PTQ.
+SmoothQuant and OmniQuant primarily act on the dynamic range of activations and weights, aiming to make low-bit quantization (typically 4–8 bits) more robust by redistributing scaling factors or applying nonlinear rescaling.
+In contrast, DBellQuant is designed specifically for near 1-bit weight quantization.
+Our learnable per-channel transform $T$ is optimized with a dedicated objective (DTMD + DTNP) that explicitly drives each channel’s weight distribution from a unimodal “single-bell” shape toward a bimodal “double-bell” shape that is theoretically more compatible with sign-based binarization.
+  • Coupled design of weight binarization and activation smoothing.
+QuaRot applies orthogonal rotations (Hadamard + permutation) to reduce correlations and outliers in weights and activations, which is highly effective for low-bit fixed-point GEMM but is not specifically tailored to the peculiarities of sign-binarized weights.
+DBellQuant, on the other hand, jointly designs (i) a learnable per-channel scaling of weights $W\odot T$ and (ii) its inverse on activations $X\odot T^{-1}$ so that:
+    1. 1.
+      the transformed weights become bimodal and thus incur smaller sign-binarization error; and
+    2. 2.
+      channels with large activation outliers are automatically assigned larger $T_{j}$, so that $T_{j}^{-1}$ suppresses those outliers when applied to activations, making 6-bit activation quantization feasible under a binary weight regime.
+  • Binary-centric loss design.
+While existing methods typically optimize reconstruction error or rely on standard calibration losses, DBellQuant’s LTDB algorithm is explicitly constructed around binary quantization.
+The Dual-Target Minimum Deviation (DTMD) and Dual-Target Normalized Proportional (DTNP) losses are defined with respect to the two binarization centers derived from the sign function, and the optimization of $T$ is directly driven by the goal of minimizing sign-binarization error.
+In other words, both the transformation and its learning objective are co-designed for binary PTQ, rather than for generic low-bit quantization.
+To summarize, although DBellQuant also employs per-channel transformations like SmoothQuant and interacts with activation outliers like QuaRot, its core contribution is conceptually different: we use a lightweight, equivalence-preserving, learnable transform to reshape the full-precision weight distributions into a bimodal form that is inherently more compatible with sign-based binarization, while the inverse transform automatically smooths activation outliers.
+This binary-centric distribution-shaping is what enables DBellQuant to achieve near 1-bit weights together with 6-bit activations under a purely post-training setting.
+A.5 Effectiveness of DTMD and DTNP Loss
+With DTMD loss, T tends to shrink toward zero, leading to a "loss hack" effect that prevents the desired transformation.
+To address this, we introduce the DTNP loss to ensure a smooth transition.
+However, using DTNP alone creates a new issue, as its convergence direction can conflict with DTMD.
+As shown in Tab.
+21, using DTNP alone drives the model toward a direction misaligned with DTMD, resulting in degraded performance.
+A.6 Hyperparameter setting in training
+Our method involves a few key hyperparameters, including the loss coefficient , the number of training epochs, and the learning rate.
+These were initially determined by validating on the LLaMA2-7B model, where we used 100 for both loss coefficients, 200 epochs, and a learning rate of 0.01.
+For the smooth parameter $\epsilon$ , we followed the strategy of SmoothQuant, testing values of 0.75, 0.80, 0.85, and 0.90.
+The differences in performance were marginal, with 0.85 yielding the best results, and thus selected as the default.
+The stopping criterion is based on LTDB loss mentioned in Section.
+3.3.
+Importantly, we found that these hyperparameters generalize well across different architectures and scales—including both the LLaMA and OPT model families—without requiring re-tuning.
+This demonstrates that our method is robust and not sensitive to hyperparameter choices.
+A.7 Visualization of the transformation of weight distributions across different layers
+The visualization of weights distribution before and after our proposed method DBellQuant is shown in Fig. 6.
 Figure 6: Top: Visualization of single-bell weights distribution from different blocks of different layers before applying DBellQuant.
 Bottom: Visualization of dual-bell weights distribution from different blocks of different layers after applying DBellQuant.
-Activation Bit-width Comparisons
-We compare the results across different models with different activation bit-widths shown in Fig.
-7.
-Specifically, we observe that when activations are quantized to 8 bits, the model’s perplexity remains nearly unchanged and, in some cases, even decreases.
-This demonstrates that our approach alleviates the challenges associated with activation quantization.
-Furthermore, we found that even under lower-bit quantization, such as 6-bit, our model is able to largely maintain its original performance, with only a minimal increase in perplexity.
-It is noteworthy that for large-scale models such as LLaMA-2 13B and 70B, the perplexity degradation is almost negligible, highlighting the effectiveness and efficiency of our approach in models with substantial parameters.
-Method Avtivation Bits Block Size WikiText2 $\downarrow$ C4 $\downarrow$
-BiLLM 16 64 20.12 24.46
-DBellQuant 8 64 13.67 16.99
-DBellQuant 6 64 15.65 19.71
-[…]
-DBellQuant 8 128 18.65 23.02
-DBellQuant 6 128 21.69 25.91
-[…]
-DBellQuant 8 256 22.34 23.37
-DBellQuant 6 256 24.68 28.52
-Table 3: Performance of different block size.
+A.8 Detailed application of Transformation
+To preserve computational consistency and avoid introducing additional parameters, we apply the transformation $T$ between the LayerNorm and the q/k/v projection layers, as well as between the LayerNorm and the up/gate projection layers.
+This placement mirrors that of SmoothQuant, ensuring the transformed inputs are compatible with quantization while maintaining the model’s architecture.
+For o_projo\_proj, although it is theoretically possible to insert a transformation $T$ between v_projv\_proj and o_projo\_proj to complete the symmetry, we found that this leads to conflicting transformations on v_projv\_proj, degrading performance.
+Therefore, we do not apply any transformation to o_projo\_proj in practice.
+Similarly, for down_projdown\_proj, the presence of activation functions (e.g., GeLU) makes it non-trivial to apply transformations without disrupting the computation graph or introducing inconsistencies.
+As a result, we do not insert transformations in this part either.
+But we need to clarify that both o_projo\_proj and down_projdown\_proj’s weight are binarized in DBellQuant.
+To clarify, for these two layers, o_projo\_proj and down_projdown\_proj, we did not learn a T-transformation to modify their weights.
+However, we still applied the original binarization in BiLLM to process these two layers, making our approach a w1a6 method.
+A.9 Performance of different activation bit-widths
+The performance of different activation bit-widths across different models is shown in Fig. 7.
 Figure 7: Performance of different activation bit-widths.
-Method Loss Function Avtivation Bits WikiText2 $\downarrow$ C4 $\downarrow$
+A.10 Results about 4-bit activation on Wikitext2 and C4.
+With the help of NVFP4 format on activation quantization, we have conducted experiments on Llama-2-7B to test the perplexity on Wikitext2 and C4 dataset, with nearly 1-bit weight and 4-bit activation.
+Results can be seen in Tab 9.
+Method Bit-width Wikitext2 C4
+fp16 16 5.47 7.26
+BiLLM 16 32.48 39.38
+DBellQuant 8 18.65 23.80
+DBellQuant 6 21.69 30.24
+DBellQuant 4 23.04 29.90
+Table 9: 4-bit activation quantization results.
+A.11 Results about training of different methods
+Detailed training time and perplexity of different methods across various models are shown in Tab. 10.
+Table 10: Training Time and Performance Comparison
+OPT-1.3B OPT-2.3B OPT-6.7B LLaMA-1-7B LLaMA-2-7B
+mins ppl mins ppl mins ppl mins ppl mins ppl
+BiLLM 7 69.97 13 49.55 34 35.36 35 35.04 37 32.48
+ARB-LLMX 88 21.81
+DBellQuant 9 43.42 17 31.47 45 18.89 47 15.34 49 17.91
+A.12 Results about bit-width, model size and perplexity of different methods
+Detailed bit-width, model size and perplexity of different methods across various models are shown in Tab. 11, Tab. 12 and Tab. 13.
+Table 11: Bit-width, Model Size and Performance Comparison for OPT-6.7B
+Methods bit-width model size ppl
+FP16 16 12.5GB 10.86
+RTN 2 2.15GB 2e4
+GPTQ 2 2.15GB 50.19
+PB-LLM 1.7 1.95GB 105.16
+BiLLM 1.11 1.84GB 35.36
+DBellQuant 1.18 1.89GB 18.89
+Table 12: Bit-width, Model Size and Performance Comparison for LLaMA-1-7B
+Methods bit-width model size ppl
+FP16 16 13.5GB 5.68
+RTN 2 2.25GB 1e5
+GPTQ 2 2.25GB 152.31
+PB-LLM 1.7 2.05GB 102.36
+BiLLM 1.09 1.97GB 35.04
+DBellQuant 1.15 2.02GB 15.34
+Table 13: Bit-width, Model Size and Performance Comparison for LLaMA-2-7B
+Methods bit-width model size ppl
+FP16 16 13.5GB 5.47
+RTN 2 2.31GB 1e5
+GPTQ 2 2.31GB 60.45
+PB-LLM 1.7 2.08GB 69.20
+BiLLM 1.08 1.98GB 32.48
+DBellQuant 1.15 2.04GB 17.91
+A.13 Results about Computation Speedup
+To quantify computational speedup, we follow the benchmarking strategy of ARB-LLM and utilize the BitBLAS codebase, which supports mixed-precision GEMM operations for low-bit weights.
+We benchmark the latency (in milliseconds) of linear layers in LLaMA2-7B using input sequences of length 2048.
+The results are shown in Tab.
+14.
+Here, we use INT8 format for activation quantization.
+Our findings are as follows:
+1)Weight binarization significantly reduces inference latency compared to FP16 models, thanks to faster memory access and bitwise computation.
+2)DBellQuant supports 8-bit and even 4-bit activation using NVFP4 format quantization, enabling efficient inference.
+This not only accelerates runtime but also yields better perplexity compared to prior binary weight methods like BiLLM and ARB-LLMX.
+Table 14: Computation Speed Comparison
+Model 4096×4096 4096×11008 11008×4096
+FP16 0.79463 1.73942 1.82653
+BiLLM 0.36842 0.38744 0.43906
+ARB-LLMX 0.33180 0.35539 0.36792
+DBellQuant 0.27694 0.30085 0.31860
+A.14 Comparison and Combination with Rotation-based Methods
+Rotation-based PTQ methods, such as QuaRot with Hadamard-based transforms, effectively reduce outliers in weights and activations.
+We compared low-bit quantization using these methods alone and in combination with our approach.
+As shown in Tab.17 and Tab.18, rotation-based PTQ methods experience significant performance degradation when weight precision is reduced below 4-bit.
+Therefore, we focus on comparisons with state-of-the-art methods for near-1-bit weight quantization.
+Additionally, we evaluated the combination of DBellQuant and Hadamard-based transforms, such as QuaRot, on the Llama-2-7B model, but this combination performed worse than DBellQuant alone, as shown in Tab.
+20.
+A possible explanation is that DBellQuant transforms the original unimodal distribution into a binarization-friendly bimodal distribution, while Hadamard-based transforms disrupt this structure, negatively impacting quantization performance.
+Method Loss Function Activation Bits WikiText2 $\downarrow$ C4 $\downarrow$
 DBellQuant L2 8 18.90 24.11
 DBellQuant L1 8 18.65 23.02
-[…]
+DBellQuant L2 6 22.26 26.41
 DBellQuant L1 6 21.69 25.91
-(a) Performance of different loss function.
+Table 15: Performance of different loss function
 Method $\epsilon$ WikiText2 $\downarrow$ C4 $\downarrow$
@@
 DBellQuant 0.9 18.29 22.52
-(b) Performance of different $\epsilon$
-Table 4: Ablation study on LLaMA-2-7B,results are measured by perplexity.
-5 Conclusion
-In this work, we propose DBellQuant, an efficient PTQ method that enables simultaneous weight binarization and activation quantization for LLMs.
-By analyzing weight distributions conducive to binarization, we design a novel and efficient Learnable Transformation for Dual-Bell algorithm.
-Specifically, we introduce two customized loss functions and an early stopping mechanism to achieve the desired dual-bell transformation.
-Furthermore, we analyze and demonstrate how this transformation improves activation quantization, allowing us to achieve near 1-bit weight compression and 6-bit activation quantization with minimal performance degradation—marking the first time such results have been achieved under a PTQ framework.
-Experiments on open-source LLM families show that DBellQuant significantly advances the performance frontier of SOTA binary PTQ methods.
-Appendix A Appendix
-A.1 Limitations
-Currently, our work only supports quantizing activations to 6 bits.
-When attempting to quantize activations to lower bit-widths, the model collapses, resulting in a significant drop in performance.
-However, some recent studies have demonstrated the ability to quantize activations to 4 bits while maintaining competitive model performance.
-Inspired by these advancements, we aim to adopt similar approaches to further optimize our models.
-Specifically, we seek to not only binarize weights but also quantize activations to even lower bit-widths, enabling easier deployment and faster computation.
-A.2 Broader Impacts
-Our work demonstrates the feasibility of simultaneously binarizing weights and quantizing activations to 6 bits in LLMs while maintaining competitive performance.
-This approach significantly reduces the computational and memory overhead associated with LLM deployment, making them more accessible for resource-constrained environments.
-By enabling efficient inference, our method contributes to the democratization of advanced AI technologies, reducing the environmental impact of large-scale model deployment.
-Furthermore, it opens new avenues for research in ultra-low-bit quantization, fostering innovation in model efficiency and scalability.
-A.3 Proof for Theorem. 1
+Table 16: Performance of different $\epsilon$
+Method 2 bit 1.1 bit
+QuaRot+RTN inf inf
+QuaRot+GPTQ 22.07 Inf
+DBellQuant - 17.91
+Table 17: Llama-2-7B results.
+Method 2 bit 1.1 bit
+QuaRot+RTN inf inf
+QuaRot+GPTQ 10.41 Inf
+DBellQuant - 12.79
+Table 18: Llama-2-13B results.
+Method Activation Bits Block Size WikiText2 $\downarrow$ C4 $\downarrow$
+BiLLM 16 64 20.12 24.46
+DBellQuant 8 64 13.67 16.99
+DBellQuant 6 64 15.65 19.71
+BiLLM 16 128 32.48 40.52
+DBellQuant 8 128 18.65 23.02
+DBellQuant 6 128 21.69 25.91
+BiLLM 16 256 43.69 43.21
+DBellQuant 8 256 22.34 23.37
+DBellQuant 6 256 24.68 28.52
+Table 19: Performance of different block size.
+Method wikitext2 C4
+DBellQuant 17.91 21.83
+DBellQuant+QuaRot 62.37 76.59
+Table 20: Combination results
+Loss Function wikitext2 C4
+DTMD+DTNP 17.91 21.83
+only DTNP 28.74 31.42
+Table 21: Loss Function results
+Method Model mins $\downarrow$ WikiText2 $\downarrow$
+BiLLM LLaMA-1-7B 35 35.04
+[…]
+DBellQuant LLaMA-1-7B 47 15.34
+BiLLM LLaMA-2-7B 37 32.48
+DBellQuant LLaMA-2-7B 49 17.91
+Table 22: Training time comparison.
+Method Bit-width $\downarrow$ Model Size $\downarrow$ Perplexity $\downarrow$
+FP16 16 13.5GB 5.47
+RTN 2 2.31GB 1e5
+GPTQ 2 2.31GB 60.45
+PB-LLM 1.7 2.08GB 69.20
+BiLLM 1.08 1.98GB 32.48
+DBellQuant 1.15 2.04GB 17.91
+Table 23: Bit-width and model size comparison.
+A.15 Perplexity on C4 dataset
+As shown in Tab.24, we compare the perplexity of the OPT and LLaMA families across different model sizes on C4 dataset.
+[…]
+Method Activation Bits OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-1-7B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
+Full Precision 16 16.07 14.34 12.71 7.34 07.26 06.73 5.71
+RTN 16 9999.56 23492.89 9617.07 194607.78 115058.76 46250.21 314504.09
+GPTQ 16 6364.65 6703.36 5576.82 186229.5 67954.04 19303.51 13036.32
+PB-LLM 16 168.12 222.15 104.78 76.63 80.69 184.67 NAN
+BiLLM 16 64.14 44.77 42.13 46.96 39.38 25.87 17.30
+[…]
+DBellQuant 16 42.57 32.89 21.78 17.60 21.83 15.14 9.49
+BiLLM 8 74.56 61.99 40.91 47.13 40.91 21.45 17.72
+DBellQuant 8 44.60 32.52 21.56 18.16 23.80 15.56 9.61
+BiLLM 6 7348 13445.21 63.41 61.65 63.41 37.66 19.43
+DBellQuant 6 57.14 45.24 23.12 19.80 30.24 17.84 10.12
+A.16 Visualization of distribution of activation before and after DBellQuant
+In this section, we present the changes in the distribution of activation values before and after applying DBellQuant as shown in Fig.8.
+It is evident that the extreme values in the activation have been significantly reduced by a factor of 5 to 10; for instance, the maximum value decreases from approximately 3 to around 0.4.
+Previous studies have highlighted that one of the primary challenges in low-bit quantization of activations lies in the presence of large outliers, which expand the activation range and, consequently, amplify quantization errors.
+By applying DBellQuant, the activation range is effectively compressed from [-3, 3] to [-0.4, 0.4], dramatically alleviating the difficulty of quantization.
+This reduction in range establishes highly favorable conditions for further exploration of lower-bit quantization, such as 8-bit or even 6-bit implementations.
+Figure 8: Visualization results of distribution of activation before and after DBellQuant across different blocks.
+A.17 Analysis of relative error of the activation
+To further validate our method, we conducted experiments demonstrating its effectiveness in reducing the relative error of activations.
+Specifically, we randomly sampled 128 data points from the C4 dataset and extracted the $q_{\text{proj}}$ inputs from the LLaMA2-7B model using both BiLLM (without applying the inverse transform on activations) and our proposed DBellQuant (with the inverse transform applied).
+We evaluated the relative error using two metrics: the Z-score, as you suggested, and the relative deviation error. The formulas are as follows:
+  • Z-score:
+    $Z=\frac{x-\mu}{\sigma}$
+    where $x$ is the value, $\mu$ is the mean, and $\sigma$ is the standard deviation.
+  • Relative Deviation Error:
+    $\text{Relative Deviation Error}=\frac{x-u}{u}$
+    where $u$ is the mean value of the corresponding row.
+These metrics allowed us to comprehensively assess the relative error and validate the robustness of our proposed method. The computation results are shown in Tab. 25.
+Z-score Relative Deviation Error
+BiLLM 0.1584 33.78
+DBellQuant 0.1401 30.72
+Table 25: Comparison of Z-score and Relative Deviation Error between BiLLM and DBellQuant.
+Our method demonstrates its effectiveness not only in reducing the absolute error of activations but also in minimizing outliers from a relative distance perspective, as verified by DBellQuant.
+A.18 Algorithm
+The detailed algorithm is shown in Algorithm. 1.
+Algorithm 1 Learnable Transformation for Dual-Bell Transformation (LTDB)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+A.19 Proof for Theorem. 1
 Proof.
@@
 ∎
-A.4 Analysis of the reasons doubel-bell distribution more suitable for binarization compared to a single-bell distribution.
+A.20 Analysis of the reasons double-bell distribution more suitable for binarization compared to a single-bell distribution.
 Directly proving that a dual-bell distribution is more suitable for binarization compared to a single-bell distribution can be challenging, as it requires setting numerous additional conditions.
@@
 Proof.
-A.4.1 Scaling Up Small Values Reduces Quantization Error
+A.20.1 Scaling Up Small Values Reduces Quantization Error
 Consider the scenario where the input vector is $X=[2,2,2,\dots,2]_{n}$.
-Assume the weights in a single channel, $W$, are given by $[\alpha_{1},\alpha_{2},...,\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $[\alpha_{i}\in U_{1}]$ is the set of values with very large absolute magnitudes, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $[\beta_{i}\in U_{2}]$ are values with magnitudes close to zero, satisfying $\sum_{i=1}^{n-k}\beta_{i}=0$.
+Assume the weights in a single channel, $W$, are given by $\alpha_{1},\alpha_{2},...,\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{i}\in U_{1}]$ is the set of values with very large absolute magnitudes, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\beta_{i}\in U_{2}]$ are values with magnitudes close to zero, satisfying $\sum_{i=1}^{n-k}\beta_{i}=0$.
 $W=U_{1}\cup U_{2}$.
@@
 $\begin{split}AbsMean&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|\beta_{1}-\frac{A}{n}|+|\beta_{2}-\frac{A}{n}|+\dots+|\beta_{n-k}-\frac{A}{n}|}{n}\end{split}$ (9)
-Given that $[\beta_{1},\beta_{2},\dots,\beta_{n-1}]$ are all very small in magnitude compared to $\frac{\alpha}{n}$, we can simplify the above as:
-$\begin{split}AbsMean&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-\beta_{1})+(\frac{A}{n}-\beta_{2})+\dots+(\frac{A}{n}-\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{2kA}{n})}{n}\end{split}$ (10)
+Given that $\beta_{1},\beta_{2},\dots,\beta_{n-1}]$ are all very small in magnitude compared to $\frac{\alpha}{n}$, we can simplify the above as:
+$\begin{split}AbsMean&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-\beta_{1})+(\frac{A}{n}-\beta_{2})+\dots+(\frac{A}{n}-\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{kA}{n})}{n}\end{split}$ (10)
 Here, the sum of the $\beta_{i}$ vanishes due to their zero sum constraint.
@@
 and for each $\beta_{i}$:
-$-AbsMean+M=-\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (12)
+$AbsMean+M=-\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (12)
 To reduce the quantization error associated with the small-magnitude weights, we can scale them up by a factor $m>1$, while correspondingly scaling down the input.
@@
 This demonstrates that scaling up the small weights does not affect the mean or absolute mean, but it can improve the quantization error, as we analyze next.
-Let us now analyze the quantization error. The original output is $2\alpha$. For convenience, let $N=(n-k)(-AbsMean+M)$. The quantized output is a sum of the dequantized values for all weights.
+Let us now analyze the quantization error.
+The original output is $2\alpha$.
+For convenience, let N=(nk)(AbsMean+M)N=(n-k)(-AbsMean+M).
+The quantized output is a sum of the dequantized values for all weights.
 $A>0$ Both before and after scaling, the quantization output contains the term:
@@
 This is because the quantized value is always less than the original due to the averaging effect.
-For the term $-AbsMean+M$:
+For the term AbsMean+M-AbsMean+M:
 $\begin{split}-AbsMean+M&=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}\\ &=-\frac{A}{n}+\frac{2kA}{n^{2}}\\ &=\frac{A}{n}\left(\frac{2k}{n}-1\right)\end{split}$ (19)
@@
 In summary, scaling up the small weights (and correspondingly scaling down the input) does not change the original computation, but it systematically reduces the quantization error by making the quantized output more faithful to the original.
-A.4.2 Scaling Down Large Values Reduces Quantization Error
+A.20.2 Scaling Down Large Values Reduces Quantization Error
 Now, let us consider the scenario where we scale down the large-magnitude weight.
-Let the input $X=[1,1,1,\dots,1]_{n}$, and the weights $W=[m\alpha_{1},m\alpha_{2},...,m\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are large values, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\alpha_{i}>0$, $[\beta_{1},\dots,\beta_{n-k}]$ are small, and $\sum_{i=1}^{n-k}\beta_{i}=0$.
+Let the input $X=[1,1,1,\dots,1]_{n}$, and the weights $W=[m\alpha_{1},m\alpha_{2},...,m\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are large values, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\alpha_{i}>0$, $\beta_{1},\dots,\beta_{n-k}]$ are small, and $\sum_{i=1}^{n-k}\beta_{i}=0$.
 The output is:
@@
 and for each $\beta_{i}$:
-$-AbsMean+M=-\frac{2m(A-\frac{A}{n})}{n}+\frac{mA}{n}=\frac{-mA+\frac{2mkA}{n}}{n}$ (26)
+$AbsMean+M=-\frac{2m(A-\frac{A}{n})}{n}+\frac{mA}{n}=\frac{-mA+\frac{2mkA}{n}}{n}$ (26)
 To scale down the large value $m\alpha$, we divide it by $m$ ($m>1$) and multiply the corresponding input element by $m$.
@@
 and for each $\beta_{i}$:
-$-AbsMean_{new}+M_{new}=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{-A+\frac{2kA}{n}}{n}$ (32)
+$AbsMean_{new}+M_{new}=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{-A+\frac{2kA}{n}}{n}$ (32)
 Let us now examine the quantization error. The original output is $m\alpha$. The quantization output before scaling is:
@@
 $\begin{split}&km(AbsMean_{new}+M_{new})+(n-k)(-AbsMean_{new}+M_{new})\\ &=km\frac{3A-\frac{2kA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &=k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\end{split}$ (34)
-Because $A>0$, for $n>100$ ,$\alpha>0$ and $k\ll n$ it holds that:
+Because $A>0$, for $n>100$ $\alpha>0$ and $k\ll n$ it holds that:
 $0<k\frac{3mA-\frac{2mkA}{n}}{n}<mA$ (35)
@@
 If $\alpha_{i}<0$, the proof is similar.
-A.5 Perplexity on C4 dataset
-As shown in Tab.5, we compare the perplexity of the OPT and LLaMA families across different model sizes on C4 dataset.
-[…]
-Method Activation Bits OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-1-7B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
-Full Precision 16 16.07 14.34 12.71 7.34 07.26 06.73 5.71
-RTN 16 $9$.56 23492.89 9617.07 194607.78 115058.76 46250.21 314504.09
-GPTQ 16 6364.65 6703.36 5576.82 186229.5 67954.04 19303.51 13036.32
-PB-LLM 16 168.12 222.15 104.78 76.63 80.69 184.67 NAN
-BiLLM 16 64.14 44.77 42.13 46.96 39.38 25.87 17.30
-[…]
-[…]
-BiLLM 8 74.56 61.99 40.91 47.13 40.91 21.45 17.72
-[…]
-BiLLM 6 7348 13445.21 63.41 61.65 63.41 37.66 19.43
-[…]
-A.6 Visualization of distribution of activation before and after DBellQuant
-In this section, we present the changes in the distribution of activation values before and after applying DBellQuant as shown in Fig.8.
-It is evident that the extreme values in the activation have been significantly reduced by a factor of 5 to 10; for instance, the maximum value decreases from approximately 3 to around 0.4.
-Previous studies have highlighted that one of the primary challenges in low-bit quantization of activations lies in the presence of large outliers, which expand the activation range and, consequently, amplify quantization errors.
-By applying DBellQuant, the activation range is effectively compressed from [-3, 3] to [-0.4, 0.4], dramatically alleviating the difficulty of quantization.
-This reduction in range establishes highly favorable conditions for further exploration of lower-bit quantization, such as 8-bit or even 6-bit implementations.
-Figure 8: Visualization results of distribution of activation before and after DBellQuant across different blocks.